A Complete Bibliography of Publications in the Journal of Open Source Software

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

13 September 2018 Version 1.00

Title word cross-reference

1 [Hug18]. 3 [Har17a, Hir18, MPFRA17, OGBC18]. ^{3} [Rei18].

-D [Hir18].

11 [Ano18a]. 14 [MB17b].

2.0 [GGKG16].

3 [CEL^{+}18, Fas18, HB16].

abundance [BG18]. Access [Ode18, VB16, Raa16]. ACE [VHM17]. ACEInhibPKPD [VHM17]. acronym [WHLM17]. Active [CHG^{+}16, Hon16]. Active-subspaces [CHG^{+}16]. add [Moe18]. add-On [Moe18]. adding [BG18]. adjoint [RFF^{+}17]. adjoint-enabled [RFF^{+}17]. ADS [GV17]. Advanced [Cuc16, RS17a]. aerosol [TCR18]. aesop [MDW18]. against [PTK16]. agent [VCD18, Den18]. agent-based [VCD18]. agnostic [SM18]. AIR [VRT16]. aka [BC16]. algebra [SC16]. algebraic [JCS17]. Algorithm [Fro16, Kea17, Por18, SS18, Tau16]. Algorithms [How18, BGR17, VBM^{+}18]. alignments [CR17]. allow [BM18]. Alphasense [HTT18]. America [MB18]. AmgX [CB17]. AmgXWrapper [CB17]. Analyse [Kle18]. Analyses [HP17, BWB^{+}17, Spi18]. Analysis [BMR^{+}16, CH17b, DC17, Gov18, HU17, ISMA18, JL16, OZW18, RS17a, RMF18, SR16, YS18, CN17, Cuc16, GJS18, Gri17, Hug18, Mäk16, NMGB17, OCT^{+}17, RRD^{+}18, SVM^{+}17, SM18, SAC^{+}17, TPAP18, WWDS18, WPK^{+}18, Wil18a, WD18, And18b, CH17a]. analytical [Ren17]. analyze [BMOF17, MB17a]. analyzing [HFF^{+}17]. AncesTrim [NSL17]. animal [MG18a]. Animation [LGH^{+}18, vdH18]. annotated [PST^{+}16]. anodes [HF16]. anomaly [All18]. anonymization [KM17b]. Anscombe [Har17b]. anywhere [TPAP18]. API [Bal16, BG17, Gin18, Smi17]. App [BSHG16, VHM17]. Application [YS18, ZKM^{+}16, GVM18]. Applications [ZKM^{+}16, And18a, CMKB18, LBČ17]. Approximate [FV18]. Approximation [MHSG18]. approxposterior [FV18]. ARC [MDW18]. architectures [CMKB18]. Arctic [Hop17]. Armadillo [SC16]. Aronnax [DR18]. array [Fas18, Lat17]. array split [Lat17]. arrgh [Kor18]. Artificial [How17]. artists [Har17a]. assembled [TSH18]. assemblies [PTK16, PST^{+}16]. assembly [APT^{+}18]. Assertive [Tho18]. Assessment [Joy17, Mut17a, Mut17b]. associated [Zek18]. Astronomical [WHLM17, HJPB17]. astrovisualization [BAB^{+}17]. atmospheres [dVBCMC17]. atmospheric [IESdF18, TCR18]. ATNF [Pit18]. Augmentation [BSH17]. Augmentor [BSH17]. Australian [SPPP17]. automated [ML16, TCR18, War16]. Automatic [GV17, WHLM17, FI18, Tan18]. Automation [OZW18]. autoplotly [Tan18]. Avoid [Att16]. aware [CR17, Gri17].

B [MPFRA17]. B-mode [MPFRA17]. BAMnostic [SM18]. Bank [Van18]. Based [ABGF17, BSHG16, Moe18, RMF18, BBM18, DDG17, HJC^{+}18, Ira18, Kea17, KM17b, KMB17, LH18, MHA17, RS17b, Rei18, SC16, SAB^{+}16, SB17, VCD18]. basic [Mor17a]. Basis [DTR18a]. batch [LBS17]. batchtools [LBS17]. Bathymetric [Hop17]. BATMAN [RRD^{+}18]. Bayesian [Gov18, JG17, TSS16]. BayesianNetwork [Gov18]. beautiful [BC16]. BeeNestABM [VCD18]. Behavioral [SD18]. benthic [Che18]. Best [Sco17]. between [CB17, DOS17, Sta17]. Bézier [Her17]. BibLaTeX [McL17]. bibliography [GV17]. BibTeX [McL17]. bikedata [PE17]. BIL [BM18]. Billiard [Dat17]. Binary [Boi18, Lar18]. Bioengineering [Moe18]. bíogo [KPA17, KA17, KSMA17]. bíogo/hts [KPA17]. bíogo/ncbi [KA17]. bioinformatics [KSMA17]. biological [SAC^{+}17]. Biomarker [HCH17]. Biomolecular [HHSP17, CCFB16]. BioPandas [Ras17]. biotmle [HCH17]. Bitstream [Boi18]. bittrex [Kan17]. blast [PTK16]. body [Gra18]. bomrang [SPPP17]. bor [SD18]. both [HB16]. Boundaries [MB18]. Boundary [CFCB17, CCFB16, CMKB18, KMB17]. Boundary-integral [CFCB17]. box [TCR18]. box-model [TCR18]. Brazilian [Ale17]. Brief [FM18]. Brightway [Mut17a]. Broadening [JGR^{+}18]. browse [BBM18]. Bruker [FSV^{+}17]. Bruker2nifti [FSV^{+}17]. bsym [Mor17a]. build [Soc18b]. building [HHV17, LBČ17, ML16, VBM^{+}18]. built [KPV^{+}17]. Bumblebees [How17, VCD18]. Bureau [SPPP17].

C [Ano18a, BM18, MB17b, ODP18, SC16, SC17, SBL^{+}17, Smi18]. C# [Lau17]. caching [SNR18]. Calculate [FN17, HF16, TSS16]. calculating [Gre16]. Calculation [Wag18, MR18b]. Calculations [OZW18, GJS18, Gra18]. Calliope [PP18]. camel [Toc18]. Canada [LA18]. Canadian [Alb17]. cancer [OCT^{+}17]. Capture [Wil18b, MG18a]. capture-recapture [MG18a]. Carbohydrate [Hon16]. Carbohydrate-Active [Hon16]. Cardiac [Ira18, RFF^{+}17]. care [SPSH^{+}17]. carl [LCP16]. Carlo [Gas18b, Mor17b]. cartography [GL16]. cartopy [Hop17]. catalogue [Pit18]. categorical [MSSH18]. Category [MSSH18]. CAZy [Hon16]. CAZy-parser [Hon16]. cbcbeat [RFF^{+}17]. Cell [HAHR18]. census [Smi17]. centers [HJC^{+}18]. central [SS18]. ChainConsumer [Hin16]. Change [LA18]. chaos [Dat18]. Chart [Hop17]. Charts [Anh18]. ChebTools [Ano18a]. Chebyshev [Ano18a]. CheckQC [ÅBD18]. checkr [Tho18]. chemicals [GVM18]. chemistry [Dah18a, TCR18]. ChemPy [Dah18a]. CheSweet [GVM18]. chloroExtractor [APT^{+}18]. chloroplast [APT^{+}18]. chronovise [RMF18]. Cine [dVBCMC17]. CircleCI [Soc18b]. CL [Spa17]. class [RS17b]. classes [SC17]. classification [Sta17, TSH18]. client [CH17a, HJC^{+}18, Kan17, Lig17]. Climate [HCM^{+}18, LA18, GWM18, Ros18, WHG17]. Climatology [Spa17]. ClimDown [HCM^{+}18]. CLIMLAB [Ros18]. CLIP [ABVF18]. CLOOPSy [DDG17]. cloud [NEGZG18]. cluster [HHV17]. Clustering [Van16, MHA17]. Coalition [BB18]. coalitions [BB18]. Cobbler [War16]. coco [Sta17]. code [Huc18b, KMB17, WZ18]. CodeMeta [Boe17b]. codes [ABEY18, RRD^{+}18]. Coding [NSB17]. cold [SS18]. cold-spells [SS18]. Collation [Tau16]. collected [MG18a]. Collection [BSHG16, ML16]. collocation [MvdB18]. Combinator [Siu17]. Combining [Wil18b]. cometary [dVBCMC17]. Command [GJS18, Soc18c]. Command-line [GJS18]. Commission [Lee16]. comorbidity [Gas18a, Gas18a]. Comparison [vH17b, FR18]. complete [TPAP18]. complex [NSL17, SK17]. ComPlot [vH17b]. Comprehensive [DC17]. Computation [AB17, ABGF17]. Computational [SP17, RFF^{+}17]. computations [NS18]. compute [dB18]. computer [RRD^{+}18]. Computing [WCP18, Cár17b, Gas18a, Lan18, Ras18, WWDS18, WFA^{+}17]. Concurrent [Gol16]. Conditional [Sco17]. Confidence [PRH17]. configuration [Raa16]. confusion [HJHZ18]. Confusograms [BC16]. Connectivity [RS17a]. consensus [CR17]. Consistent [MBK^{+}18]. containerized [HHV17]. Containers [Soc18a]. Containershare [Soc18b]. Contemporary [MB18]. context [BGR17, Wak16b]. contigs [PHSK17]. continuous [WPK^{+}18]. contraction [SG18]. contrib [MSSH18]. contribute [BBM18]. control [ÅBD18, CMEM^{+}17, JA16, MB17b, ODP18]. conversion [MPFRA17, McF16]. conversions [Hir18]. convert [ABEY18, GZT^{+}18]. converter [FSV^{+}17, Sta17]. converting [Sta17]. Convex [YK17]. Convolutional [Mor17c]. coordinate [Hir18]. Copernicus [DDG17]. CoreRobotics [ODP18]. corner.py [FM16]. Correlation [ALB^{+}16, MR18b]. correspondence [RS17b]. cottoncandy [NEGZG18]. country [ABEY18, Sta17]. countrycode [ABEY18]. covariance [BC16]. cOver [DDG17]. Crawford [BF17]. Create [GL16, OGBC18]. Creating [MD17, Joy17, Mur18]. Creation [GGKG16]. Criteria [Bos17]. Critical [SPSH^{+}17]. Cross [CH18, ODP18]. cross-language [ODP18]. cross-platform [ODP18]. Cross-Validation [CH18]. crOwdsourcing [DDG17]. CRU [Spa17]. Crypto [Kan17]. Crypto-Currency [Kan17]. CSS [LeB18a]. CTL [ALB^{+}16]. cuIBM [KMB17]. Currency [Kan17]. curve [BF17]. Curves [Her17]. Cycle [Joy17, Mut17a, Mut17b]. cycles [Hug18]. Cycling [OSS18]. Cylc [OSS18].

D [Har17a, Hir18, Hug18, MPFRA17, OGBC18]. DAE [MKT^{+}18]. Daily [SHN17]. DART [LGH^{+}18]. Data [Alb17, BSHG16, Boi18, CCAM18, DC17, GHF^{+}17, HHSP17, Hug16, JL16, KMG^{+}17, Kle18, Lee16, LIK18, Lüd18a, Lüd18b, Ode18, PC18, SR16, SD18, SPPP17, SHN17, Tie17, VB16, Van18, Vit17, Wro18, APT^{+}18, Bil18, CN17, Dir18, FM18, Gin18, GZT^{+}18, Har17b, HFF^{+}17, JA16, LA18, MB17a, MG18a, MSSH18, MZ18, NMGB17, Ras18, RC18, SPSH^{+}17, Smi17, WWDS18, WPK^{+}18, Zek18]. Data-Efficient [CCAM18]. Database [Hon16, FI18, VRT16]. databases [GV17]. DataFrames [Ras17]. dataset [RS17b]. datasets [HJPB17]. datastructures [Dir18]. Decomposition [DTR18b]. Deep [Arn17, DC17, Fay17]. defined [Dah18b]. defining [RS17b]. Deflagrations [HH16]. DEFRA [VRT16]. degree [Lev16]. democracies [BB18]. Density [LIK18, JNM18, MHA17, Woj17]. deploy [Soc18b]. Design [MH18]. detection [All18, Ano18b, Mor17c, SS18]. Detonations [HH16]. different [Sta17]. differential [Kar18]. Dimensional [Hon17, Lat17]. dimensionality [Mad16]. dimensions [Dat17]. direct [MvdB18]. Directory [Att16]. Discovery [HCH17, Vit17]. Discrete [And18a, JEC18, vdH18]. disease [TR17]. Display [Hug16]. dissertation [Ale17]. Distance [Dro18]. distributed [CMKB18, SNR18]. distributed-memory [CMKB18]. distribution [VCD18]. Distributions [FV18]. dit [JEC18]. Django [HDL17]. dms2dfe [DC17]. DNA [BI16, War16]. Document [GGKG16]. Documents [LeB18a]. DoSOCS [GGKG16]. Download [LA18, vH17a]. Downloading [PC18]. Downscaling [HCM^{+}18]. draft [War16]. drake [Lan18]. draw [RC18]. DSL [Bal16]. dust [Gre18]. dustmaps [Gre18]. Dynamic [CM17, DTR18b, LGH^{+}18, BG18, CM18]. dynamical [Mar17, Dat17]. DynamicalBilliards.jl [Dat17]. DynamicalSystems.jl [Dat18]. Dynamics [HHSP17, HP17, Dat18, PW17, dB18].

ease [TPAP18]. easier [CN17]. easy [Dat17, Har17a, NEGZG18, RRD^{+}18, vH17a, DTR18a]. easy-to-use [Dat17, Har17a]. Eating [FN17]. Echelle [MDW18]. Eclipse [PLL^{+}16]. ECNet [KM17a]. ecology [Che18]. Ecopath [Kea17]. ecopath matlab [Kea17]. EDA [KPV^{+}17]. EDAM [BBM18]. edarf [JL16]. Education [BA18]. EFAshiny [YS18]. efd [Gri17]. effective [Wha18, SP17]. Effective-Quadratures [SP17]. Effects [Lüd18a]. Efficient [CCAM18, Mak18, McF16]. effmass [Wha18]. EggNOG [FI18]. egtplot [MWS18]. einsum [SG18]. einsum-like [SG18]. eixport [IESdF18]. eLabFTW [CMP17]. Elastic [GHF^{+}17]. Electrograms [RGZ^{+}18]. electron [Woj17]. Electrophysiology [Ira18, RFF^{+}17]. electrostatics [CFCB17, CCFB16]. Elegant [Inn18]. Elektra [Raa16]. element [LH18]. Elements [Spa17, CCFB16, TSH18]. elliptical [Gri17]. emass [Por18]. embedding [GF18]. EMBL [PST^{+}16]. emissions [IESdF18]. Empirical [CM17]. ENA [BG17]. enabled [RFF^{+}17]. ENASearch [BG17]. Encoders [MSSH18]. encoding [MSSH18]. Encounter [Car17a]. EndoMineR [Zek18]. endoscopic [Zek18]. Energy [BWB^{+}17, HHM18, PP18]. Engine [OSS18, SB17]. Engineering [SP17]. enhanced [ABVF18]. ensemble [VWDB16]. Ensight4Matlab [SBL^{+}17]. EnSight(R) [SBL^{+}17]. ensuring [RS17b]. entity [KM17b]. entity-based [KM17b]. ENVI [BM18]. ENVI-BIL [BM18]. Environment [LA18]. environments [HHV17, Hir18]. enZYmes [Hon16]. EQ [SP17]. equation [Kar18]. equations [Dah18b, DDJ^{+}17, WZ18]. Equivalent [MH18]. Errors [AB17]. estimate [Zag18]. estimates [Lev16, SR18]. Estimation [Rov17, DRS16, JNM18]. Euclidean [Van16]. evaluate [vH17b]. evaluating [ML16]. evaluator [Huc18a]. Event [Hal17, NSB17, vdH18]. Evolutionary [MWS18]. Exchange [Lee16, Kan17]. excitation [dVBCMC17]. Exclusion [Fin18]. executing [WFA^{+}17]. execution [Spi18, SB17]. Exoplanet [Ano18b]. expansions [Ano18a]. Expectation [SC17]. Expectation-Maximisation [SC17]. expensive [RRD^{+}18]. Experiment [Soc18a]. Experimental [RGZ^{+}18]. Experiments [MD17, ABVF18]. Exploratory [YS18, CN17, JL16]. Explore [Spa17]. Exponential [Wro18, Lev16]. export [IESdF18]. expression [Har17b]. expressions [SG18]. extendable [Dat17]. extension [MPFRA17]. extensions [Ras18]. Extract [Alb17, Hon16]. Extraction [RS17a, APT^{+}18, Zek18]. Extractor [Bar16]. Extraordinarily [Did17]. ezknitr [Att16]. EZyRB [DTR18a].

f [Rei18]. facilitating [Spi18]. Factor [YS18]. Factory [Soc18a]. Fall [MH18]. Family [Wro18]. Fast [AEAP18, BS16, MBK^{+}18, NS18, PHSK17, CEL^{+}18]. Feature [RS17a, Mad16]. Fetch [SPPP17]. Feynman [Fro18]. fgivenx [Han18]. fib [Ira18]. fib-tf [Ira18]. Fiducialized [RGZ^{+}18]. field [DRS16]. fields [Mur18]. file [Did17, Lar18, NPP17]. Files [EW16, LeB18b, Huc18b, SVM^{+}17, SBL^{+}17]. filesstrings [NPP17]. fill [GV17]. filltex [GV17]. filtering [BG18]. Financial [Lee16]. Finch [BG18]. finishing [War16]. finite [LH18]. finreportr [Lee16]. firm [BMOF17]. Fitting [Gol16, MG18a]. Flexible [CCAM18, GHF^{+}17, CEL^{+}18, Por18]. Flow [DDJ^{+}17]. fluorescence [dVBCMC17]. flusight [TR17]. Flux [Inn18]. FNFT [WCP18]. focused [Woj17]. food [Kea17]. forecasts [TR17]. foreground [Joy17]. forensic [And18a]. Forest [BGR17]. Forests [PRH17, JL16]. format [Fas18, FSV^{+}17, LA18, SBL^{+}17]. Formatting [LeB18a]. Fourier [Gri17, WCP18]. Frames [Lüd18a, Tie17]. Framework [CH18, Den18, GHF^{+}17, RGZ^{+}18, BAB^{+}17, Che18, CMEM^{+}17, HHV17, Mar17, Mut17a, OCT^{+}17, PP18, Raa16, RMF18, RFF^{+}17, SAB^{+}16, Som17, Tay18]. framwork [Rei18]. free [HFF^{+}17, LCP16, MG18a, PHSK17]. free-ranging [MG18a]. FRIEDA [GHF^{+}17]. Friendly [YS18]. friends [OGBC18]. fuel [KM17a]. Full [DDJ^{+}17, Hal17]. FullSWOF [DDJ^{+}17]. fully [Joy17]. Function [Rov17]. Functional [Gol16, Wro18, Han18]. Functions [CN17, Lüd18b, BF17, May17, MR18b, WPK^{+}18]. fuse [VWDB16]. fuzzy [TSS16].

G [Rei18]. Gala [PW17]. galactic [PW17]. Galore [JGR^{+}18]. Games [MWS18]. Gamma [MG18a]. Gas [Mor17b]. Gaussian [CCAM18, SC17]. gem [BGR17, GF18]. gene [AEAP18, Har17b]. GeneNetwork [SAB^{+}16]. General [Lar18, Kar18]. Generalized [CH18]. generated [HF16]. Generating [Boe17b, ZKM^{+}16, OCT^{+}17]. generation [MPFRA17, Tan18]. Generator [Bos18, TCR18, WZ18, Hag17]. generators [Lau17]. generic [CMEM^{+}17]. genetics [And18a, SAB^{+}16]. geneXplain [SKW17]. geneXplainR [SKW17]. Genome [TSH18, APT^{+}18, OCT^{+}17, War18]. Genome-wide [TSH18]. genomepy [vH17a]. genomes [War16, vH17a]. Genomic [Hic16, BGR17, BG18, SM18].

Genotify [AEAP18]. genotype [ALB^{+}16]. Geocoding [Hal17]. geometrically [Lev16]. geometrically-weighted [Lev16]. Geometry [Moe18]. Geoparsing [Hal17]. geospace [Hir18]. Geospatial [CH17a]. geostatistics [Hof18]. GeoStats.jl [Hof18]. getCRUCLdata [Spa17]. GFF3toEMBL [PST^{+}16]. ggeffects [Lüd18a]. Giant [BC16]. GIBBON [Moe18]. Gillespie.jl [Fro16]. GIS [Cuc16, Mut17b]. Git [JA16]. Git-RDM [JA16]. Global [SHN17, Rei18]. glycan [GVM18]. gmm diag [SC17]. gmm full [SC17]. Go [BWB^{+}17, KPA17, KA17, KSMA17, Kor18]. Go-HEP [BWB^{+}17]. Gold [MG18b, SBL^{+}17]. Golo [PLL^{+}16]. Gompertz [MG18a]. good [HB16]. Government [SPPP17]. GPU [CFCB17, KMB17]. GPU-based [KMB17]. GPUs [CCFB16]. gradient [Rei18]. gradient-based [Rei18]. Grafoscopio [Cár17b]. Gramm [Mor18]. grammar [Mor18]. Graph [Bos18, Hag17, Boe17a, GF18, Lev16]. graph-theoretic [Boe17a]. grapherator [Bos18]. graphical [KPV^{+}17]. graphics [Mor18]. graphs [Wak16a]. gravitational [WPK^{+}18]. gravitational-wave [WPK^{+}18]. Greek [Van18]. groundwater [Rei18]. GSODR [SHN17]. GWAS [Tur18]. gwdegree [Lev16].

h5preserve [Toc18]. h5py [Toc18]. Habfuzz [TSS16]. habitat [TSS16]. Handle [GZT^{+}18]. handling [Did17, KPA17]. hdbscan [MHA17]. hddtools [Vit17]. Healthy [FN17]. heatmaps [SK17]. heatwaveR [SS18]. heatwaves [SS18]. Hebbian [MS16]. hebbRNN [MS16]. Hector [WHG17]. HEI [FN17, FN17]. Height [MH18]. Helper [Her17, Soc18c]. HelpMe [Soc18c]. HEP [BWB^{+}17]. Herfindahl [Wag18]. hhi [Wag18]. Hierarchical [MHA17]. High [BWB^{+}17, CCAM18, Hof18, Hon17, KPA17, KSMA17, Lan18, MKT^{+}18, WFA^{+}17]. High-Dimensional [Hon17]. High-performance [CCAM18, Hof18, KSMA17, Lan18]. Higher [Her17]. highlightHTML [LeB18a]. Hirschman [Wag18]. Histogram [RS17a]. Histogram-weighted [RS17a]. Historical [MB18]. Hits [Sco17]. hmis [MB17a]. homology [WWDS18]. HPC [HJC^{+}18]. HSImage [BM18]. htmlwidget [Coe18]. hts [KPA17]. humanleague [Smi18]. Humans [Boi18]. hydraulic [Ren17, TSS16]. Hydrological [Van18, VWDB16, Vit17]. Hydrometric [Alb17]. hydroscoper [Van18]. Hyperparameter [MHH^{+}16]. hyperspectral [BM18]. HyPhy [Spi18]. Hytool [Ren17].

I/O [NPP18]. IBCAO py [Hop17]. idealised [DR18]. identification [MvdB18, TSH18]. identifiers [Huc18b]. IDESolver [Kar18]. iheatmapr [SK17]. ijtiff [NPP18]. Illumina [ÅBD18]. Image [BSH17, Moe18]. Image-Based [Moe18]. ImageJ [NPP18]. Images [FSV^{+}17, ISMA18, Soc17, WHLM17, BM18, Mor17c]. Imaging [Woo18, ZB17, MPFRA17]. ImagingReso [ZB17]. iml [MCB18]. immersed [CMKB18, KMB17]. immersed-boundary [CMKB18]. implement [BGR17]. Implementation [Fro18, BF17, Gri17, Kea17, Lig17, Por18, Sco17, Tau16]. implementing [BMOF17]. Import [Kle18, McL17]. improved [Sco17]. Improvement [Anh18]. Improving [Lev16]. including [Den18, SR18]. indel [CR17]. indel-aware [CR17]. independent [FR18]. Index [FN17, Wag18]. Inelastic [LIK18]. infectious [TR17]. inference [ALB^{+}16, LCP16, TSS16]. inflationary [MR18b]. Information [Dro18, HP17, Gra18, Hon16, JEC18, LBČ17, Van18].

Information-Theoretic [HP17]. infrared [dVBCMC17]. Infrastructure [Den18, ZKM^{+}16]. inhibition [VHM17]. initio [GJS18]. INSPIRE [GV17]. inspired [Toc18, VBM^{+}18]. instream [TSS16]. instrumental [Zag18]. integral [CFCB17]. Integrate [GL16]. integrated [Mäk16]. integration [Dah18c]. Integrator [Hon17]. integrity [RS17b]. integro [Kar18]. integro-differential [Kar18]. Intelligent [GHF^{+}17]. Interact [VRT16]. interacting [BG17]. interaction [BM18]. Interactive [Gov18, SK17, Joy17, OCT^{+}17, Ros18, Som17, Tan18, TR17, Wak16a]. Interface [Arn17, CB17, Coe18, DOS17, Gre18, HJPB17, Kor18, MKT^{+}18, Pit18, PWFM17, SKW17, Van18, WZ18, WHG17]. interfaces [KA17, Smi17, Smi18]. interferometeric [HJPB17]. Intermittently [RGZ^{+}18]. International [Hop17]. interoperable [LBČ17]. Interpretable [MCB18]. interpretation [Lev16, Ren17]. interstellar [Gre18]. Intervals [PRH17]. Intro.js [Gan16]. Introducing [KM17b]. IP [KPV^{+}17]. IP-XACT [KPV^{+}17]. isopycnal [DR18]. isotropic [Mur18]. iterative [BHA18]. ivporbit [Zag18].

Java [Den18]. Javascript [CH17a, Por18]. js [Por18]. js-emass [Por18]. jstor [Kle18]. Julia [BH18, Dat17, Dat18, Fro16, Hof18, Inn18, Lau17, MR18a, Pas17]. Jump [MH18].

Kactus2 [KPV^{+}17]. Keras [Arn17]. kerasR [Arn17]. kernel [JNM18]. key [RS17b]. key-based [RS17b]. khmer [SAC^{+}17]. kima [Ano18b]. Kindel [CR17]. knitr [Att16]. Kraljic [BMOF17]. KraljicMatrix [BMOF17].

laboratory [CMP17]. Labour [Ode18]. labs [CMP17]. Land [DDG17]. landscapes [Cuc16]. Language [MBK^{+}18, JCS17, KPA17, KA17, KSMA17, Mäk16, ODP18]. languages [Mäk16]. Laplace [And18a]. Large [KM17a, SNR18]. large-scale [SNR18]. LAST [Sco17]. LaTex [GV17]. Lattice [Mor17b]. Lattice-Gas [Mor17b]. lattice mc [Mor17b]. Launcher [WFA^{+}17]. lawn [CH17a]. Layer [Den18]. LCA [Joy17, CM18]. Lcopt [Joy17]. learn [MSSH18]. Learning [Arn17, BSH17, Fay17, HCH17, MHH^{+}16, MS16, MCB18, BF17, CEL^{+}18, Inn18, KM17a, RS17b, Ras18]. learningCurve [BF17]. LearnSAT [BA18]. level [MKT^{+}18]. libqcpp [MB17b]. libraries [BWB^{+}17]. Library [BSH17, CH17a, CHG^{+}16, CCAM18, Fay17, Gan16, ISMA18, Tor18, WCP18, Bar16, BG17, BI16, BM18, CB17, Coe18, CEL^{+}18, Dat18, Gre16, HJHZ18, HU17, HB16, Hop17, LBČ17, May17, MB17b, ODP18, RC18, SC16, SH17, WPK^{+}18, Arn17]. Life [How17, Joy17, Mut17a, Mut17b]. light [SB17]. Lightweight [Fay17, AEAP18]. like [SG18]. likelihood [LCP16]. likelihood-free [LCP16]. Limarka [Ale17]. Limbo [CCAM18]. Line [Soc18c, dVBCMC17, GJS18]. Lineage [And18b]. Linear [Gol16, Dah18b, SC16]. lists [WD18]. literate [Cár17b]. Loci [ALB^{+}16]. log [JNM18]. log-transformed [JNM18]. logic [TSS16]. logKDE [JNM18]. long [War16]. lookup [AEAP18]. LSPR [CFCB17].

M [Rei18]. M-f [Rei18]. M2M [MG18b]. Machine [BSH17, MHH^{+}16, MCB18, CEL^{+}18, Inn18, KM17a, RS17b, Ras18]. macromolecular [Woj17]. made [RRD^{+}18]. MAGICC [GWM18]. Magnetic [FSV^{+}17]. maintaining [RS17b]. making [Ale17, CN17]. malan [And18b]. MAle [And18b]. Manage [McL17, RC18]. Management [GHF^{+}17, JA16]. Manhattan [Tur18]. Manifold [MHSG18]. Manipulate [You18, GZT^{+}18].

Manipulation [Hic16, NPP17]. Manuscripts [PC18]. many [Gra18]. many-body [Gra18]. Mapelia [OGBC18]. mapped [Fas18]. Mapping [Wak16b, ALB^{+}16, DDG17]. Maps [GL16, Gre18, OGBC18]. Marginal [Lüd18a]. Markdown [Ale17, LeB18a]. Market [Ode18]. Masks2Metrics [MG18b]. mass [Wha18]. MassMine [VB16]. Matching [KB18]. MatchPy [KB18]. MatDL [Fay17]. mathematical [JCS17, MR18a]. MATLAB [FR18, Fay17, Ren17, VHM17, Kea17, MG18b, Mor18]. Matlab-based [Kea17]. MATLAB/Octave [FR18]. MATLAB(R) [SBL^{+}17]. matplotlib [Hop17, RC18]. Matrices [SD18, FM16]. matrix [HJHZ18, SH17, BMOF17]. Maximisation [SC17]. MCAPL [Den18]. MCMC [YK17, You18]. MCMCvis [You18]. mcMST [Bos17]. MCycle [Hug18]. MDEntropy [HP17]. Measurement [RMF18]. Measurement-Based [RMF18]. Medical [ISMA18, Zek18]. memory [CMKB18, Fas18]. memory-mapped [Fas18]. Metadata [Boe17b]. Meteorological [Van18]. Meteorology [SPPP17]. method [CMKB18, DTR18a, KMB17]. Methods [Gol16, GF18]. metrics [Gre16]. micompm [FR18]. MicroBenthos [Che18]. microbial [Che18]. microcomputers [HTT18]. microcontrollers [HTT18]. microcontrollers/microcomputers [HTT18]. microframework [VBM^{+}18]. MicroPEM [SVM^{+}17]. microPIECE [ABVF18]. microRNA [ABVF18]. microsynthesis [Smi18]. Microtiter [Hug16]. MinHash [BI16]. MinHashing [BG18]. Minimum [Bos17, Van16]. Mining [SR16]. missing [Bil18]. Missingno [Bil18]. mixed [Wak16b]. mixture [And18a, SC17]. mlpack [CEL^{+}18]. MLxtend [Ras18]. mmappickle [Fas18]. modalities [MPFRA17]. Mode [DTR18b, MPFRA17]. Model [Gol16, And18a, DR18, Fro18, GWM18, MG18a, TCR18, VHM17, VCD18, WHG17, Zag18, vH17b]. Modeling [CM17, Gov18, ISMA18, Che18, CCAM18, HHM18, JCS17, Mar17, MZ18, Ros18]. Modelling [Car17a, VWDB16, PP18, Rei18, Tay18]. Models [Lüd18a, IESdF18, Joy17, Lev16, ML16, OGBC18, SC17]. Modular [Bos18, Dat17]. Modulated [MS16]. Module [MD17, Mor17b, Fas18, Mor17a]. moldable [Cár17b]. Molecular [HP17, Ras17, dB18]. molten [LH18]. Moltres [LH18]. Monte [Gas18b, Mor17b]. Mordecai [Hal17]. Morphometrics [MG18b]. Mosquitto [Lig17]. mpnum [SH17]. MQTT [Lig17]. MSMExplorer [HHSP17]. mst clustering [Van16]. Multi [Bos17, Bos18, Lat17, BB18, PP18, SC17]. Multi-Criteria [Bos17]. Multi-dimensional [Lat17]. multi-party [BB18]. multi-scale [PP18]. Multi-Step [Bos18]. multi-threaded [SC17]. Multiclass [HJHZ18]. Multilocus [PTK16]. Multiphonon [LIK18]. Multiple [Mad16, BH18, Mäk16]. multiple-scattering [BH18]. multivariate [FR18]. Mutational [DC17].

N2 [HTT18]. Named [KM17b]. names [ABEY18, Sta17]. National [Van18, SR18]. Natural [MBK^{+}18]. nature [VBM^{+}18]. nature-inspired [VBM^{+}18]. ncbi [KA17]. Negative [YK17]. nested [CN17, WD18]. nests [VCD18]. NETANOS [KM17b]. Network [Gov18, Mor17c, ALB^{+}16, Gre16, WZ18]. Networks [MS16, RS17a, Boe17a]. Neural [MS16, Mor17c]. neurophysiology [NMGB17]. Neuropsydia.py [MD17]. Neuroscience [RS17a]. Neutron [LIK18, ZB17]. next [MPFRA17]. next-generation [MPFRA17]. NGS [CDLC17]. NiaPy [VBM^{+}18]. Nifti [FSV^{+}17]. Nimbus [BGR17]. Node.js [KM17b]. noisy [May17]. Noisyopt [May17]. Nomis [Ode18]. nomisr [Ode18]. nomisweb [Smi17]. Non [Wil18b, YK17, Dah18b]. non-linear [Dah18b]. Non-Negative [YK17]. Non-Probability [Wil18b]. Nonlinear [WCP18, Dat18]. nonsense [Huc18a]. Nostril [Huc18a]. notebook [CMP17]. novo [PTK16]. nse [AB17]. nuclear [WZ18]. nucleotide [CR17]. Numerical [AB17, Dah18c]. numerically [Dah18b]. NumPy [DOS17, EW16, RC18, Fas18]. NVIDIA [CB17].

O [NPP18]. object [Har17a, ODP18]. object-oriented [ODP18]. Objects [Her17]. Observation [MDW18, SD18]. observations [FR18]. ocean [DR18, vH17b, Hop17]. OctApps [WPK^{+}18]. Octave [FR18, WPK^{+}18]. ODE [Dah18c, MKT^{+}18]. ODES [MKT^{+}18]. OnDemand [HJC^{+}18]. ontology [BBM18]. opc [HTT18]. OPC-N2 [HTT18]. OPEM [HAHR18]. Open [Fin18, HAHR18, HJC^{+}18, KM17b, Soc17, BAB^{+}17, CM18, CMP17, HU17, Mut17a, Ren17, VHM17, VCD18, Soc18b]. open-source [BAB^{+}17, HU17, VHM17, VCD18]. OpenCPU [Kor18]. OpenEPSim [Fin18]. OpenSpace [BAB^{+}17]. OpenStreetMap [Boe17a]. operate [HTT18]. opt einsum [SG18]. Optim [MR18a]. Optimization [CCAM18, JG17, MHH^{+}16, Mir18, JCS17, MR18a, MvdB18]. optimizing [BH18, May17, SG18]. Optlang [JCS17]. opty [MvdB18]. Order [Her17, SG18]. organisation [Dir18]. Oriented [Mak18, ODP18, Ros18]. origami [CH18]. OS-agnostic [SM18]. osmdata [PLSR17]. OSMnx [Boe17a]. Osprey [MHH^{+}16]. other [HTT18]. Output [You18, SVM^{+}17].

Overland [DDJ^{+}17]. OWL [Bal16].

p5.js [Har17a]. Package

[Mak18, MWS18, Rov17, SD18, Wag18, ABEY18, BMOF17, Boe17a, BHA18, Car17a, Dah18a, Dat17, Dir18, FM18, Gas18a, GF18, Gra18, Han18, HHM18, Hug18, IESdF18, JEC18, Lan18, Lau18, MSSH18, MR18a, MCB18, MZ18, MR18b, Mur18, NMGB17, NPP17, NPP18, NEGZG18, PW17, RG18, SVM^{+}17, Smi18, SG18, Spi18, Sta17, Tan18, Tho18, Tur18, VWDB16, Wha18, Wil18a, WD18, Zag18, dB18]. Packages [Boe17b]. Packer [RCS^{+}16]. Pain [Att16]. Pandarus [Mut17b]. pandas [Ras17]. parallel [PWFM17]. parameter [BC16, MvdB18]. parameterised [Joy17]. parameters [Raa16]. ParaVision [FSV^{+}17]. Parser [Siu17, Lar18, Hon16]. parsing [Spi18]. Particle [Mir18]. ParticleScattering [BH18]. partitioning [Lat17]. party [BB18]. Pathfinder [Den18]. pathology [Zek18]. Pattern [KB18]. PDF [LeB18b]. pdfsearch [LeB18b]. pedigrees [NSL17]. PEM [HAHR18]. performance [CCAM18, Hof18, KSMA17, Lan18]. periodic [GJS18]. persistent [WWDS18]. PetIBM [CMKB18]. PETRARCH2 [NSB17]. PETSc [CB17]. PFEIFER [RGZ^{+}18]. pharmacodynamic [VHM17]. pharmacokinetic [VHM17]. pharmacokinetic/pharmacodynamic [VHM17]. phenology [Tay18]. phenotype [ALB^{+}16]. Philentropy [Dro18]. Phonetic [How18]. Phonon [LIK18]. photoelectron [JGR^{+}18]. Phylen [FI18]. Phylogemetric [Gre16]. phylogenetic [FI18, Gre16, NS18, WD18]. phylogenomic [PHSK17]. phylogram [WD18]. phyphy [Spi18]. Physics [BWB^{+}17]. pi [HTT18]. picker [Har17a]. pickle [Fas18]. Pipeline [BMR^{+}16, TPAP18, WHLM17, ABVF18, Lan18, SB17, WWDS18, MDW18]. pipelines [CDLC17]. Pipengine [SB17]. plant [Tay18]. plater [Hug16]. Plates [Hug16]. Platform [DDG17, ODP18, SKW17]. Plotrr [CN17]. plots [BC16, Tur18]. Plotter [vH17b]. Plotting [OZW18, GJS18, Han18, HFF^{+}17, Mor18]. plug [Cuc16]. plug-in [Cuc16]. plugin [JA16]. polaron [Fro18]. PolaronMobility.jl [Fro18]. Pollution [VRT16, RC18]. Polynomials [SP17]. Polytopes [YK17]. pools [PWFM17]. popeye [DRS16]. popular [HTT18]. population [DRS16, SR18]. populations [MG18a]. portal [HJC^{+}18]. portfolio [BMOF17]. Portfolios [ABGF17]. possible [Ale17]. Posterior [FV18, Han18]. Powder [LIK18]. power [Hug18, Mur18]. powerbox [Mur18]. predict [GVM18]. prediction [KM17a]. predictive [MZ18]. Preparing [PST^{+}16]. Preprocessing [RGZ^{+}18]. Principles [SR16]. Prism [Mad16]. Prison [FM18]. prisonbrief [FM18]. Probabilistic [RMF18]. probabilities [BB18]. Probability [Wil18b]. probit [Zag18]. Problem [Bos17, HH16]. problems [BH18]. Process [Fin18, Ros18, SBL^{+}17]. process-oriented [Ros18]. Processes [CCAM18]. Processing [BMR^{+}16, Tor18, PWFM17, SPSH^{+}17]. processor [MZ18]. prodest [Rov17]. product [SH17]. Production [Rov17, BF17]. Program [NSB17, HF16]. Programming [Tho18, Bal16]. Projection [MHSG18]. projections [SR18]. projects [KM17a, SNR18]. property [KM17a]. protocol [Lig17]. Providing [Ras18, NPP18]. pseudo [PHSK17]. pseudo-phylogenomic [PHSK17]. psrqpy [Pit18]. psycho [Mak18]. Psychological [Mak18]. PsychoPhysioCollector [BSHG16]. Psychophysiological [BMR^{+}16, BSHG16]. PsychoPhysioPipeline [BMR^{+}16]. psyplot [Som17]. Published [PC18].

Publishing [Mak18]. Publishing-Oriented [Mak18]. PubMLST [PTK16]. pulsar [Pit18]. purchasing [BMOF17]. purpose [Kar18]. pvlib [HHM18]. py [HTT18]. py-opc [HTT18]. PyBox [TCR18]. PyCM [HJHZ18]. PyDMD [DTR18b]. PyGBe [CFCB17, CCFB16]. PyGBe-LSPR [CFCB17]. pyGPGO [JG17]. pygtc [BC16]. pyhector [WHG17]. Pymagicc [GWM18]. PyMap3D [Hir18]. pyneqsys [Dah18b]. pynucastro [WZ18]. pyodesys [Dah18c]. pyPhenology [Tay18]. Pyret [NMGB17]. Pyrgg [Hag17]. pysrim [OZW18]. PySwarms [Mir18]. python [Gra18, HJPB17, HHM18, MB17a, NEGZG18, Pit18, Smi17, Tay18, Ano18a, BG17, Boe17a, BHA18, BM18, CFCB17, CH17b, CHG^{+}16, CCFB16, Dah18a, Dah18c, DTR18b, Fas18, FV18, FM16, GWM18, GZT^{+}18, GF18, Gre18, Gre16, Hag17, HJHZ18, Han18, HU17, HFF^{+}17, HB16, Huc18a, Hug18, ISMA18, JEC18, JG17, KB18, MD17, Mar17, May17, McF16, Mir18, MWS18, Mor17b, MR18b, Mur18, NMGB17, PRH17, PW17, PWFM17, RS17b, RG18, Ras18, Ros18, RC18, Sco17, Smi18, SG18, Spi18, Sta17, SH17, Tau16, VBM^{+}18, WHG17, vdH18, HHM18]. PyTransport [MR18b]. pyuca [Tau16]. PyUnfold [BHA18]. pyuvdata [HJPB17].

Q [Tur18]. Q-Q [Tur18]. qicharts2 [Anh18]. qqman [Tur18]. Quadratures [SP17]. Quail [HFF^{+}17]. Quality [Anh18, ÅBD18, MB17b]. qualtRics [Gin18]. Quantification [Dro18, TSH18]. QUantitative [Woo18]. Quantum [Cuc16, Gra18]. quaternions [RG18]. queries [GV17]. querying [Pit18]. Questionnaires [MD17]. Quick [ÅBD18]. quimb [Gra18]. QUIT [Woo18].

R [Anh18, AB17, ABGF17, ABEY18, Arn17, BMOF17, BS16, Boe17b, Car17a, CH17a, Coe18, CH18, Dir18, Dro18, EW16, FM18, Gas18a, GL16, Hic16, HCM^{+}18, How18, IESdF18, Kan17, Kor18, Lan18, LBS17, Lau18, LeB18a, McL17, MCB18, MZ18, NPP17, NPP18, Rov17, SVM^{+}17, SK17, SNR18, SR16, Siu17, Smi17, Smi18, SD18, SPPP17, SHN17, Spa17, SKW17, Tan18, Tho18, Tur18, Van18, VWDB16, WWDS18, Wil18a, WD18, Zag18]. R3D2 [HH16]. radial [Ano18b]. RAILS [War16]. Ramble [Siu17]. Random [BGR17, Car17a, Hag17, JL16, PRH17, Lev16]. ranging [MG18a]. raspberry [HTT18]. rate [McF16]. rates [WZ18]. ray [HF16]. RcppCNPy [EW16]. rdefra [VRT16]. RDM [JA16]. reaction [WZ18]. Reactive [HH16]. reactors [LH18]. Read [EW16, Hug16, SBL^{+}17]. Read-Write [EW16]. reading [HB16]. reads [PHSK17, Wak16b]. recall [HFF^{+}17]. Recapture [Wil18b, MG18a]. receptive [DRS16]. Reciprocal [Sco17]. Recognition [Tor18]. reconstruction [FI18]. Recordings [RGZ^{+}18]. Records [SD18]. Recurrent [MS16]. Reduced [DTR18a]. Reduction [WHLM17, Mad16]. reference [PHSK17]. reference-free [PHSK17]. References [McL17]. RefManageR [McL17]. regionalized [Mut17b]. register [Wro18]. Registration [Wro18]. Registry [Soc17, Soc18b]. Regression [Lüd18a, Mad16]. regularization [Mad16]. Relativistic [HH16]. release [SAC^{+}17]. remBoot [Car17a]. Remote [HDL17]. reper [TSH18]. repetitive [TSH18]. Reports [Pas17, OCT^{+}17, Zek18]. Representation [Hic16, SH17]. Representations [Lau18]. Reproducibility [RCS^{+}16, Lan18]. Reproducible [Soc18a, Cár17b, HHV17, SNR18, TPAP18]. ReproZip [RCS^{+}16]. resampy [McF16]. Research [BMR^{+}16, BSHG16, ZKM^{+}16, Cár17b, CMP17, JA16, Mir18]. Resonance [FSV^{+}17, ZB17]. results [Gas18b, Tan18, Tur18]. ResumableFunctions [Lau17]. Retrieval [KMG^{+}17]. retrieve [Gin18]. Retriever [KMG^{+}17]. returns [FM18]. reusable [BBM18]. Reward [MS16]. Reward-Modulated [MS16]. Riemann [HH16]. rintrojs [Gan16]. RISE [Wil18a]. Risk [ABGF17]. Risk-Based [ABGF17]. RiskPortfolios [ABGF17]. RNA [Har17b, TPAP18]. RNA-seq [Har17b, TPAP18]. RNAsik [TPAP18]. robot [ODP18]. Robotics [LGH^{+}18]. Robust [GHF^{+}17]. ROOT [DOS17]. root numpy [DOS17]. ROS [CMEM^{+}17]. ros control [CMEM^{+}17]. rowan [RG18]. rsimsum [Gas18b]. RSMTool [ML16]. RTI [SVM^{+}17]. rtimicropem [SVM^{+}17]. Ruby [BGR17]. rucrdtw [BS16]. Rule [MS16]. runs [ÅBD18, TPAP18].

safe [NS18]. SaffronTree [PHSK17]. salabim [vdH18]. SALib [HU17]. salt [LH18]. Sample [Wil18b, McF16]. Sampling [YK17]. SAT [BA18]. Scaffolding [War16]. Scala [Bal16]. scale [KM17a, PP18, SNR18]. scan [All18, MPFRA17]. Scanning [DC17]. scanstatistics [All18]. Scattering [LIK18, BH18]. Scatterplot [FM16]. schemes [Sta17]. schwimmbad [PWFM17]. Science [KM17b, Mak18, Ras18]. Scientific [Kle18, Pas17, NEGZG18, Ras18]. scikit [MSSH18]. scikit-learn-contrib [MSSH18]. Scores [FN17, Wag18, Gas18a]. scoring [ML16]. Scowl [Bal16]. screen [Mor17c]. Search [LeB18b, BS16]. Securities [Lee16]. selection [BGR17, Mad16]. Sensitivity [HU17]. Sentinel [DDG17]. Sentinel-based [DDG17]. SEP [Bar16]. seq [Har17b, TPAP18]. ’Sequana’ [CDLC17]. sequence [CR17, KPA17, MB17b, PTK16, SM18, SAC^{+}17]. sequences [War16]. sequencing [ÅBD18]. Series [Lau18, BS16]. server [Kor18, Lig17]. services [KA17]. Set [CDLC17]. Setting [Wil18b]. Shallow [DDJ^{+}17]. Shallow-Water [DDJ^{+}17]. sharp [Lau17]. SHGYield [AM17]. shifts [GVM18]. Shiny [YS18]. shmlast [Sco17]. shotgun [APT^{+}18]. sigma.js [Coe18]. sigmajs [Coe18]. simple [CMEM^{+}17, Did17, GWM18, KSMA17, WHG17, WFA^{+}17]. simpleCache [SNR18]. simulating [Mar17]. Simulation [Fin18, Fro16, HAHR18, How17, Gas18b, JGR^{+}18, LH18, vdH18]. simulations [TCR18, dB18, vH17b]. Simulator [Ira18]. SimuPy [Mar17]. Singularity [Soc17]. sizing [Hug18]. sjmisc [Lüd18b]. skedm [CM17]. Skeleton [ZKM^{+}16]. sketching [BI16]. Ski [MH18]. skijumpdesign [MH18]. Slicer [MPFRA17]. SlicerITKUltrasound [MPFRA17]. Smartphone [BSHG16]. Smartphone-Based [BSHG16]. Snakemake [CDLC17]. Software [MvdB18, WCP18, CM18, Dat18, HHV17, SAC^{+}17]. solar [HHM18]. solutions [Ren17]. Solve [Dah18b]. Solver [BA18, CFCB17, HH16, Kar18]. solvers [MKT^{+}18]. Solving [BH18]. sound [MZ18]. Source [Bar16, HAHR18, Soc17, Soc18b, BAB^{+}17, CM18, CMP17, HU17, Huc18b, Mut17a, Ren17, VHM17, VCD18]. sourmash [BI16]. space [All18]. space-time [All18]. Spanning [Bos17, Van16]. spatial [Gri17, Gri17]. spatial-aware [Gri17]. spatial-efd [Gri17]. spatiotemporal [VCD18]. SPDX [GGKG16]. species [Wak16b]. Specified [MH18]. spectra [HF16, Mur18]. Spectral [CH17b]. Spectroscopic [MDW18]. spectroscopy [JGR^{+}18]. ’Spectrum’ [CH17b]. Speech [Tor18]. SpeechPy [Tor18]. speed [TPAP18]. spells [SS18]. Spiral [Huc18b]. spline [Mad16]. splitters [Huc18b]. SRIM [OZW18]. stabilizing [Har17b]. stack [Ras18]. Standard [AB17, MG18b, KPV^{+}17, Spi18]. States [LIK18, MB18]. Statistical [RRD^{+}18, Tan18]. statistically [MZ18]. statistics [All18]. Step [Bos18]. Stochastic [Fro16, dB18]. Storage [GGKG16, Dir18, NEGZG18]. store [Fas18]. Straightforward [Dah18c]. strategically [BMOF17]. Strategy [MWS18]. Streamlined [Wag18]. street [Boe17a]. string [Huc18a, NPP17]. structured [Mur18]. Structures [ISMA18, Ras17]. students [Ale17]. Studies [SP17, Gas18b]. style [Lau17]. subject [ALB^{+}16]. Submission [HDL17, PST^{+}16]. subnational [SR18]. subsequence [BS16]. subspaces [CHG^{+}16]. SuchTree [NS18]. suitability [TSS16]. suite [Bil18]. Summarise [Gas18b]. summarization [AEAP18]. Summarize [You18]. Summary [SHN17]. sumo [GJS18]. Suppdata [PC18]. Supplementary [PC18]. Support [EW16, HB16]. supporting [SVM^{+}17]. survey [Gin18]. survival [MG18a, Wak16a]. SurvivalVolume [Wak16a]. Swarm [Mir18]. symbolically [Dah18b]. symmetry [Mor17a]. synteny [War18]. Synthetic [ZKM^{+}16]. System [GGKG16, JA16, Kor18]. Systems [OSS18, Dah18c, Dah18b, Dat17, HHM18, LBS17, Mar17, PP18].

Targeted [HCH17]. Tasks [MD17].

TDAstats [WWDS18]. template [SC16]. template-based [SC16]. Temporalis [CM18]. TensorFlow [Ira18]. TensorFlow-based [Ira18]. terrestrial [Hir18]. test [Soc18b]. tests [Ren17]. Text [Hal17, Mor17c, MBK^{+}18, SR16, KM17b]. Texts [Kle18]. tf [Ira18]. Theoretic [HP17, Boe17a]. Theory [Dro18, JEC18]. thermodynamic [Hug18]. thesis [Ale17]. Thin [Toc18]. thread [NS18]. thread-safe [NS18]. threaded [SC17]. Three [MWS18]. Three-Strategy [MWS18]. threshold [Wak16a]. throughput [KPA17, WFA^{+}17].

Tidy [Alb17, Hug16, Lüd18a, SR16, FM18]. tidyhydat [Alb17]. tidynamics [dB18]. tidytext [SR16]. TIFF [NPP18]. Time [Lau18, All18, BS16]. Timing [RMF18]. tiny [dB18]. Tokenization [MBK^{+}18]. Tool [Fin18, HAHR18, KMG^{+}17, MH18, ZB17, BG18, Cár17b, DRS16, Joy17, KPV^{+}17, MB17a, Mäk16, NSL17, TSS16, WFA^{+}17]. Toolbox [Bos17, MG18b, CMKB18, FR18, HFF^{+}17, LCP16, Ren17]. Toolkit [ISMA18, LGH^{+}18, KSMA17, Lan18, Mir18, Mut17b, Ros18, SM18]. Tools [LBS17, LeB18b, Vit17, Woo18, You18, Ano18a, GJS18, LIK18, ML16, SPSH^{+}17]. topological [WWDS18]. Trait [ALB^{+}16]. trajectory [MvdB18]. Transformation [Lüd18b, Har17b]. transformed [JNM18]. transformers [MSSH18]. Transforming [SD18]. Transforms [WCP18]. Tree [Bos17, BBM18]. tree-based [BBM18]. Trees [Van16, NS18, PHSK17]. Triangle [BC16]. Triangles [Her17]. trimming [NSL17]. TSrepr [Lau18]. Tubular [ISMA18]. tungsten [HF16]. Tuples [Hic16]. Turf [CH17a]. two [Dat17]. Typical [Att16]. typing [PTK16].

U.S. [Lee16]. UglyMol [Woj17]. UK [Ode18, Smi17, VRT16]. UKCensusAPI [Smi17]. ukpopulation [SR18]. ultra [SB17]. ultrasound [MPFRA17]. UMAP [MHSG18]. unfolding [BHA18]. Unicode [Tau16]. unified [SR18]. Uniform [MHSG18, PWFM17]. United [MB18]. units [GZT^{+}18]. universal [Raa16]. unyt [GZT^{+}18]. USAboundaries [MB18]. Use [Spa17, Dat17, Har17a]. useful [Dah18a]. User [YS18]. User-Friendly [YS18]. users [NPP18]. Using [Att16, Pas17, SR16, All18, FI18, Gin18, Hop17, JL16, MvdB18, Ren17, TSS16, Tur18, War16]. utilities [Ras18]. Utility [CHG^{+}16, Soc18c]. v [Spa17]. v2.1 [SAC^{+}17]. Validation [CH18]. Variable [Lüd18b]. variables [Zag18]. variance [Har17b]. variants [SR18]. variational [Fro18]. Varistran [Har17b]. Vascular [ISMA18]. vbvs.concurrent [Gol16]. VCC [HHV17]. VCF [HB16]. VCFPy [HB16]. velocities [Ano18b]. version [JA16]. via [Van16]. viewer [Woj17]. viewshed [Cuc16]. Virtual [How17]. visdat [Tie17]. visual [CN17, Cuc16, Har17a]. Visualising [Tie17]. Visualization [Wag18, Bil18, BBM18, Coe18, Som17]. Visualizations [HHSP17, Tan18, TR17]. Visualize [You18, MB17a]. Visualizing [War18, Tur18]. visually [vH17b]. volume [Wak16a]. vtreat [MZ18].

walkr [YK17]. Water [DDJ^{+}17]. wave [WPK^{+}18]. way [Hon16, vH17a]. Weather [SHN17, LA18]. weathercan [LA18]. Weave.jl [Pas17]. web [BBM18, HJC^{+}18, Kea17, SAB^{+}16, KM17b]. web-based [HJC^{+}18, SAB^{+}16, KM17b]. web-visualization [BBM18]. WebGL [Woj17]. website [FM18]. weighted [Lev16, RS17a]. weighting [JGR^{+}18]. Whole [Tie17, APT^{+}18]. wide [TSH18]. wind [RC18]. windrose [RC18, RC18]. without [TSH18]. work [Boe17a, LBS17]. Workflow [DC17, GL16, Mak18, OSS18]. Working [Att16, Ras17, Ano18a, RG18]. workloads [WFA^{+}17]. World [FM18]. Wrapper [Gan16, GWM18, Toc18]. wrappers [ODP18]. Wright [BF17]. Write [EW16, SBL^{+}17]. writing [Ale17, HB16]. written [Dah18a, Huc18a].

X [HF16]. X-ray [HF16]. XACT [KPV^{+}17]. Xbim.Essentials [LBČ17]. Xenomapper [Wak16b]. xmatchview [War18]. xpecgen [HF16]. xphyle [Did17].

YAML [SB17]. YAML-based [SB17].

YoungTab [Rom17].

References

Ardia:2017:PNC

[AB17] David Ardia and Keven Bluteau.

nse: Computation of numerical standard errors in R. Journal of Open Source Software, 2(10):172:1–172:2, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00172.

Aaslin:2018:PCQ

[ÅBD18] Matilda Åslin, Monika Brandt, and Johan Dahlberg. CheckQC: Quick quality control of Illumina sequencing runs. Journal of Open Source Software, 3(22):556:1–556:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00556.

Arel-Bundock:2018:PCR

[ABEY18] Vincent Arel-Bundock, Nils Enevoldsen, and C. J. Yet-

man.  countrycode: An

R package to convert country names and country codes. Journal of Open Source Software, 3(28):848:1–848:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00848.

Ardia:2017:PRC

[ABGF17] David Ardia, Kris Boudt, and Jean-Philippe Gagnon-Fleury.

RiskPortfolios: Computa-

tion of risk-based portfolios in R. Journal of Open Source Software, 2(10):171:1, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00171.

Amsel:2018:PMM

[ABVF18] Daniel Amsel, André Billion, Andreas Vilcinskas, and Frank Förster. microPIECE — microRNA pipeline enhanced by CLIP experiments. Journal of Open Source Software, 3(24):616:1–616:4, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00616.

Andrews:2018:PGF

[AEAP18] Jared M. Andrews, Mohamed El-Alawi, and Jacqueline E. Payton. Genotify: Fast, lightweight gene lookup and summarization. Journal of Open Source Software, 3(28): 885:1–885:4, August 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00885.

Arends:2016:CTL

[ALB^{+}16] Danny Arends, Yang Li, Gudrun A. Brockmann, Ritsert C. Jansen, Robert W. Williams, and Pjotr Prins. Correlation trait loci (CTL) mapping: phenotype network inference subject to genotype. Journal of Open Source Software, 1(6):87:1–87:4, October 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00087.

Albers:2017:PTE

[Alb17]  Sam Albers.  tidyhydat:

Extract and tidy Canadian hydrometric data. Journal of Open Source Software, 2 (20):511:1–511:4, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00511.

Alexandre:2017:PLM

[Ale17] Eduardo S. M. Alexandre.

Limarka: making possible Brazilian students writing dissertation and thesis with Markdown. Journal of Open Source Software, 2(9):169:1, January 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00169.

Allevius:2018:PSS

[All18] Benjamin Allévius. scanstatistics:

space-time anomaly detection using scan statistics. Journal of Open Source Software, 3(25):515:1–515:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00515.

Anderson:2017:PS

[AM17] Sean M. Anderson and Bernardo S. Mendoza. SHGYield. Journal of Open Source Software, 2(14):242:1–242:2, June 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00242.

Andersen:2018:DLM

[And18a] Mikkel Meyer Andersen. Discrete Laplace mixture model with applications in forensic genetics. Journal of Open Source Software, 3(26):748:1– 748:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00748.

Andersen:2018:PMM

[And18b] Mikkel Meyer Andersen. malan: MAle Lineage ANalysis. Journal of Open Source Software, 3(25):684:1–684:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00684.

Anhoj:2018:PQQ

[Anh18]  Jacob Anhøj.  qicharts2:

Quality improvement charts for R. Journal of Open Source Software, 3(25):699:1– 699:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00699.

Anonymous:2018:PCC

[Ano18a]  Anonymous.  ChebTools:

C++11 (and Python) tools for working with Chebyshev expansions. Journal of Open Source Software, 3(22):569:1– 569:3, February 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00569.

Anonymous:2018:PKE

[Ano18b] Anonymous. kima: Exoplanet detection in radial velocities. Journal of Open Source Software, 3(26):487:1–487:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00487.

Ankenbrand:2018:PCE

[APT^{+}18] Markus J. Ankenbrand, Simon Pfaff, Niklas Terhoeven, Musga Qureischi, Maik Gündel, Clemens L. Weiß, Thomas Hackl, and Frank

Förster. chloroExtractor:

extraction and assembly of the chloroplast genome from whole genome shotgun data. Journal of Open Source Software, 3(21):464:1–464:3, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00464.

Arnold:2017:PKR

[Arn17] Taylor B. Arnold. kerasR: R interface to the Keras Deep Learning Library. Journal of Open Source Software, 2(14): 296:1, June 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00296.

Attali:2016:PEA

[Att16] Dean Attali. ezknitr: Avoid the typical working directory pain when using ‘knitr’. Journal of Open Source Software, 1(5):75:1, September 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00075.

Ben-Ari:2018:PLS

[BA18] Mordechai (Moti) Ben-Ari. LearnSAT: A SAT solver for education. Journal of Open Source Software, 3(24):639:1– 639:2, April 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00639.

Bock:2017:POO

[BAB^{+}17] Alexander Bock, Emil Axelsson, Karl Bladin, Jonathas Costa, Gene Payne, Matthew Territo, Joakim Kilby, Masha Kuznetsova, Carter Emmart, and Anders Ynnerman. OpenSpace: An open-source astrovisualization framework. Journal of Open Source Software, 2(15): 281:1, July 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00281.

Balhoff:2016:PSS

[Bal16] James P. Balhoff. Scowl: a Scala DSL for programming with the OWL API. Journal of Open Source Software, 1(1): 23:1, May 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00023.

Barbary:2016:PSS

[Bar16] Kyle Barbary. SEP: Source extractor as a library. Journal of Open Source Software, 1(6):58:1–58:2, October 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00058.

Bender:2018:PCC

[BB18] Andreas Bender and Alexander Bauer. coalitions: Coalition probabilities in multiparty democracies. Journal of Open Source Software, 3(23):606:1–606:2, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00606.

Brancotte:2018:RTB

[BBM18] Bryan Brancotte, Christophe Blanchet, and Hervé Ménager. A reusable tree-based webvisualization to browse EDAM ontology, and contribute to it. Journal of Open Source Software, 3(27):698:1–698:4, July 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00698.

Bocquet:2016:PPB

[BC16] Sebastian Bocquet and Faustin W. Carter. pygtc: beautiful parameter covariance plots (aka. giant triangle confusograms). Journal of Open Source Software, 1(6):46:1–46:3, October 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00046.

Boehmke:2017:PLI

[BF17] Bradley C. Boehmke and Ja-

                                        son K. Freels. learningCurve:

An implementation of Crawford’s and Wright’s learning curve production functions. Journal of Open Source Software, 2(13):202:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00202.

Batut:2017:PEP

[BG17] Bérénice Batut and Björn Grüning. ENASearch: A Python library for interacting with ENA’s API. Journal of Open Source Software, 2(18):418:1, October 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00418.

Bovee:2018:PFT

[BG18] Roderick Bovee and Nick Greenfield. Finch: a tool adding dynamic abundance filtering to genomic MinHashing. Journal of Open Source Software, 3(22):505:1–505:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00505.

Bazan:2017:PNR

[BGR17] Juanjo Bazán and Oscar Gonzalez-Recio. Nimbus: a Ruby gem to implement random forest algorithms in a genomic selection context. Journal of Open Source Software, 2(16):351:1–351:3, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00351.

Blankrot:2018:PPS

[BH18] Boaz Blankrot and Clemens

Heitzinger. ParticleScattering:

Solving and optimizing multiplescattering problems in Julia. Journal of Open Source Software, 3(25):691:1–691:3, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00691.

Bourbeau:2018:PPP

[BHA18] James Bourbeau and Zigfried Hampel-Arias. PyUnfold: A Python package for iterative unfolding. Journal of Open Source Software, 3(26):741:1– 741:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00741.

Brown:2016:PSL

[BI16] C. Titus Brown and Luiz Irber. sourmash: a library for MinHash sketching of DNA. Journal of Open Source Software, 1(5):27:1, September 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00027.

Bilogur:2018:PMM

[Bil18] Aleksey Bilogur. Missingno: a missing data visualization suite. Journal of Open Source Software, 3(22):547:1–547:4, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00547.

Brown:2018:PHP

[BM18] Ryan C. Brown and Joshua Moser. HSImage: A Python and C++ library to allow interaction with ENVI-BIL hyperspectral images. Journal of Open Source Software, 3(25):630:1–630:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00630.

Boehmke:2017:PKR

[BMOF17] Bradley C. Boehmke, Robert T. Montgomery, Jeffrey A. Ogden, and Jason K. Freels.

                                        KraljicMatrix: An R package

for implementing the Kraljic Matrix to strategically analyze a firm’s purchasing portfolio. Journal of Open Source Software, 2(10):170:1–170:2, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00170.

Bogutzky:2016:PPP

[BMR^{+}16] Simon Bogutzky, Phillip Marsch, Licínio Roque, Nassrin Hajinejad, and Barbara Grüter.

PsychoPhysioPipeline:  A

processing and analysis pipeline for psychophysiological research. Journal of Open Source Software, 1(5):41:1– 41:3, September 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00041.

Boeing:2017:POP

[Boe17a] Geoff Boeing. OSMnx: A Python package to work with graph-theoretic Open-StreetMap street networks. Journal of Open Source Software, 2(12):215:1–215:4, April 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00215.

Boettiger:2017:GPC

[Boe17b] Carl Boettiger. Generating CodeMeta metadata for R packages. Journal of Open Source Software, 2(19):454:1, November 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00454.

Boisgerault:2018:PBB

[Boi18] Sébastien Boisgérault. Bitstream — binary data for humans. Journal of Open Source Software, 3(21):541:1–541:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00541.

Bossek:2017:PMT

[Bos17]  Jakob Bossek.  mcMST: A

toolbox for the multi-criteria minimum spanning tree problem. Journal of Open Source Software, 2(17):374:1–374:2, September 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00374.

Bossek:2018:PGM

[Bos18] Jakob Bossek. grapherator: A modular multi-step graph generator. Journal of Open Source Software, 3(22):528:1–528:3, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00528.

Boersch-Supan:2016:PRF

[BS16] Philipp H. Boersch-Supan. rucrdtw: Fast time series subsequence search in R. Journal of Open Source Software, 1(7):100:1–100:2, November 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00100.

Bloice:2017:PAI

[BSH17] Marcus D. Bloice, Christof Stocker, and Andreas Holzinger. Augmentor: An image augmentation library for machine learning. Journal of Open Source Software, 2(19):432:1– 432:2, November 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00432.

Bogutzky:2016:PPS

[BSHG16] Simon Bogutzky, Jan Christoph Schrader, Nassrin Hajinejad, and Barbara Grüter.

PsychoPhysioCollector: A

smartphone-based data collection app for psychophysiological research. Journal of Open Source Software, 1(4): 40:1–40:3, August 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00040.

Binet:2017:PGH

[BWB^{+}17] Sebastien Binet, Bastian Wieck, David Blyth, Emmanuel Busato, Michaël Ughetto, and Peter Waller. Go-HEP: libraries for high energy physics analyses in Go. Journal of Open Source Software, 2(17):372:1–372:4, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00372.

Caravaggi:2017:PRR

[Car17a] Anthony Caravaggi. remBoot: An R package for random encounter modelling. Journal of Open Source Software, 2(10):176:1–176:3, February 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00176.

Cardenas:2017:PGM

[Cár17b] Offray Vladimir Luna Cárdenas. Grafoscopio: A moldable tool for literate computing and reproducible research. Journal of Open Source Software, 2(18):251:1–251:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00251.

Chuang:2017:PAI

[CB17] Pi-Yueh Chuang and Lorena A.

                                        Barba. AmgXWrapper: An

interface between PETSc and the NVIDIA AmgX library. Journal of Open Source Software, 2(16):280:1, August 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00280.

Cully:2018:PLF

[CCAM18] Antoine Cully, Konstantinos Chatzilygeroudis, Federico Allocati, and Jean-Baptiste Mouret. Limbo: A flexible high-performance library for Gaussian processes modeling and data-efficient optimization. Journal of Open Source Software, 3(26):545:1– 545:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00545.

Cooper:2016:PPP

[CCFB16] Christopher D. Cooper, Natalia C. Clementi, Gilbert Forsyth, and Lorena A. Barba. PyGBe: Python, GPUs and boundary elements for biomolecular electrostatics. Journal of Open Source Software, 1 (4):43:1, August 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00043.

Cokelaer:2017:PSSa

[CDLC17] Thomas Cokelaer, Dimitri Desvillechabrol, Rachel Legendre, and Mélissa Cardon. ’Sequana’: a set of Snakemake NGS pipelines. Journal of Open Source Software, 2(16):352:1–352:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00352.

Curtin:2018:PMF

[CEL^{+}18] Ryan R. Curtin, Marcus Edel, Mikhail Lozhnikov, Yannis Mentekidis, Sumedh Ghaisas, and Shangtong Zhang. mlpack 3: a fast, flexible machine learning library. Journal of Open Source Software, 3(26):

726:1–726:2, June 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00726.

Clementi:2017:PPL

[CFCB17] Natalia C. Clementi, Gilbert Forsyth, Christopher D. Cooper, and Lorena A. Barba. PyGBe-LSPR: Python and GPU boundaryintegral solver for electrostatics. Journal of Open Source Software, 2(19):306:1–306:2, November 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00306.

Chamberlain:2017:PLR

[CH17a] Scott Chamberlain and Jeffrey W. Hollister. lawn: An R client for the Turf Javascript Library for Geospatial Analysis. Journal of Open Source Software, 2(11):194:1, March 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00194.

Cokelaer:2017:PSSb

[CH17b] Thomas Cokelaer and Juer-

                                        gen Hasch.  'Spectrum':

Spectral analysis in Python. Journal of Open Source Software, 2(18):348:1–348:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00348.

Coyle:2018:POG

[CH18] Jeremy R. Coyle and Nima S.

Hejazi. origami: A generalized framework for cross-validation in R. Journal of Open Source Software, 3(21):512:1–512:3, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00512.

Chennu:2018:PMM

[Che18] Arjun Chennu. MicroBenthos:

a modeling framework for microbial benthic ecology. Journal of Open Source Software, 3(25):674:1–674:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00674.

Constantine:2016:PPA

[CHG^{+}16] Paul Constantine, Ryan Howard, Andrew Glaws, Zachary Grey, Paul Diaz, and Leslie Fletcher. Python active-subspaces utility library. Journal of Open Source Software, 1(5):79:1, September 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00079.

Cortale:2017:PSE

[CM17] Nicholas Cortale and Dylan McNamara. skedm: Empirical dynamic modeling. Journal of Open Source Software, 2(12):207:1–207:2, April 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00207.

Cardellini:2018:PTO

[CM18] Giuseppe Cardellini and Chris Mutel. Temporalis: an open source software for dynamic LCA. Journal of Open Source Software, 3(24):612:1–612:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00612.

Chitta:2017:PRG

[CMEM^{+}17] Sachin Chitta, Eitan Marder-Eppstein, Wim Meeussen, Vijay Pradeep, Adolfo Rodríguez Tsouroukdissian, Jonathan Bohren, David Coleman, Bence Magyar, Gennaro Raiola, Mathias Lüdtke, and Enrique Fernandez Perdomo. ros control: A generic and simple control framework for ROS. Journal of Open Source Software, 2(20):456:1–456:5, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00456.

Chuang:2018:PPT

[CMKB18] Pi-Yueh Chuang, Olivier Mesnard, Anush Krishnan, and Lorena A. Barba. PetIBM: toolbox and applications of the immersed-boundary method on distributed-memory architectures. Journal of Open Source Software, 3(25):558:1– 558:5, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00558.

Carpi:2017:PEO

[CMP17] Nicolas Carpi, Alexander Minges, and Matthieu Piel. eLabFTW: An open source laboratory notebook for research labs. Journal of Open Source Software, 2(12):146:1, April 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00146.

Crabtree:2017:PPF

[CN17] Charles Crabtree and Michael J.

Nelson. Plotrr: Functions for making visual exploratory data analysis with nested data easier. Journal of Open Source Software, 2(11):190:1–190:2, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00190.

Coene:2018:PSR

[Coe18] Jean-Philippe Coene. sigmajs:

An R htmlwidget interface to the sigma.js visualization library. Journal of Open Source Software, 3(28):814:1–814:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00814.

Constantinides:2017:PKI

[CR17] Bede Constantinides and David L.

Robertson. Kindel: indelaware consensus for nucleotide sequence alignments. Journal of Open Source Software, 2(15):282:1–282:2, July 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00282.

Cuckovic:2016:AVA

[Cuc16] Zoran Cuckovic. Advanced viewshed analysis: a quantum GIS plug-in for the analysis of visual landscapes. Journal of Open Source Software, 1(4):32:1–32:3, August 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00032.

Dahlgren:2018:PCP

[Dah18a] Björn Dahlgren. ChemPy: A package useful for chemistry written in Python. Journal of Open Source Software, 3(24):565:1–565:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00565.

Dahlgren:2018:PPSb

[Dah18b] Björn Dahlgren. pyneqsys: Solve symbolically defined systems of non-linear equations numerically. Journal of Open Source Software, 3(21):531:1– 531:2, January 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00531.

Dahlgren:2018:PPSa

[Dah18c] Björn Dahlgren. pyodesys:

Straightforward numerical integration of ODE systems from Python. Journal of Open Source Software, 3(21):490:1– 490:2, January 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00490.

Datseris:2017:PDJ

[Dat17] George Datseris. DynamicalBilliards.jl:

An easy-to-use, modular and extendable Julia package for Dynamical Billiard systems in two dimensions. Journal of Open Source Software, 2(19):458:1–458:4, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00458.

Datseris:2018:PDJ

[Dat18] George Datseris. DynamicalSystems.jl:

A Julia software library for chaos and nonlinear dynamics. Journal of Open Source Software, 3(23):598:1–598:5, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00598.

deBuyl:2018:PTT

[dB18] Pierre de Buyl. tidynamics:

A tiny package to compute the dynamics of stochastic and molecular simulations. Journal of Open Source Software, 3(28):877:1–877:4, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00877.

Dandage:2017:PDC

[DC17] Rohan Dandage and Kausik Chakraborty. dms2dfe: Comprehensive workflow for analysis of deep mutational scanning data. Journal of Open Source Software, 2(20):362:1–362:3, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00362.

DeVecchi:2017:PCC

[DDG17] Daniele De Vecchi, Fabio Dell’Acqua, and Daniel Aurelio Galeazzo. CLOOPSy — Copernicus Land cOver crOwdsourcing Platform for Sentinel-based mapping. Journal of Open Source Software, 2(15):291:1– 291:2, July 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00291.

Delestre:2017:PFF

[DDJ^{+}17] Olivier Delestre, Frédéric Darboux, François James, Carine Lucas, Christian Laguerre, and Stéphane Cordier. FullSWOF: Full shallow-water equations for overland flow. Journal of Open Source Software, 2(20):448:1, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00448.

Dennis:2018:MFI

[Den18]  Louise A. Dennis.  The

MCAPL framework including the Agent Infrastructure Layer and Agent Java Pathfinder. Journal of Open Source Software, 3(24):617:1–617:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00617.

Didion:2017:PXE

[Did17] John P. Didion. xphyle: Extraordinarily simple file handling. Journal of Open Source Software, 2(14):255:1, June 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00255.

Dirmeier:2018:PDR

[Dir18] Simon Dirmeier. datastructures:

An R package for organisation and storage of data. Journal of Open Source Software, 3(28):910:1–910:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00910.

Dawe:2017:PRI

[DOS17] Edmund Noel Dawe, Piti Ongmongkolkul, and Giordon Stark. root numpy: The interface between ROOT and NumPy. Journal of Open Source Software, 2(16):307:1–307:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00307.

Doddridge:2018:PAI

[DR18] Edward W. Doddridge and Alexey Radul. Aronnax: An idealised isopycnal ocean model. Journal of Open Source Software, 3(26):592:1–592:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00592.

Drost:2018:PPI

[Dro18] Hajk-Georg Drost. Philentropy: Information theory and distance quantification with R. Journal of Open Source Software, 3(26):765:1–765:4, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00765.

DeSimone:2016:PPP

[DRS16] Kevin DeSimone, Ariel Rokem, and Keith Schneider. popeye: a population receptive field estimation tool. Journal of Open Source Software, 1(8):103:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00103.

Demo:2018:PEE

[DTR18a] Nicola Demo, Marco Tezzele, and Gianluigi Rozza. EZyRB: Easy Reduced Basis method. Journal of Open Source Software, 3(24):661:1–661:3, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00661.

Demo:2018:PPP

[DTR18b] Nicola Demo, Marco Tezzele, and Gianluigi Rozza. PyDMD: Python Dynamic Mode Decomposition. Journal of Open Source Software, 3(22):530:1– 530:3, February 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00530.

deVal-Borro:2017:PCL

[dVBCMC17] Miguel de Val-Borro, Martin A. Cordiner, Stefanie N. Milam, and Steven B. Charnley. Cine: Line excitation by infrared fluorescence in cometary atmospheres. Journal of Open Source Software, 2(11):182:1–182:2, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00182.

Eddelbuettel:2016:PRR

[EW16] Dirk Eddelbuettel and Wush Wu. RcppCNPy: Read-write support for NumPy files in R. Journal of Open Source Software, 1(5):55:1, September 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00055.

Fasnacht:2018:PMP

[Fas18] Laurent Fasnacht. mmappickle:

Python 3 module to store memory-mapped numpy array in pickle format. Journal of Open Source Software, 3(26):651:1–651:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00651.

Fayek:2017:PML

[Fay17] Haytham M. Fayek. MatDL: A lightweight deep learning library in MATLAB. Journal of Open Source Software, 2(19):413:1–413:2, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00413.

Ferres:2018:PPA

[FI18] Ignacio Ferrés and Gregorio Iraola. Phylen: automatic phylogenetic reconstruction using the EggNOG database. Journal of Open Source Software, 3(25):593:1–593:3, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00593.

Finn:2018:POO

[Fin18] Caley Finn. OpenEPSim: An open exclusion process simulation tool. Journal of Open Source Software, 3(24): 650:1–650:2, April 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00650.

Foreman-Mackey:2016:PCP

[FM16] Daniel Foreman-Mackey. corner.py:

Scatterplot matrices in Python. Journal of Open Source Software, 1(2):24:1–24:2, June 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00024.

Freire:2018:PPR

[FM18] Danilo Freire and Robert Myles McDonnell. prisonbrief: An R package that returns tidy data from the World Prison Brief website. Journal of Open Source Software, 3(22):361:1, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00361.

Folsom:2017:PHC

[FN17] Tim Folsom and V. P. Nagraj. hei: Calculate healthy eating index (HEI) scores. Journal of Open Source Software, 2(18):417:1, October 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00417.

Fachada:2018:PMM

[FR18] Nuno Fachada and Agostinho C.

Rosa. micompm: A MATLAB/ Octave toolbox for multivariate independent comparison of observations. Journal of Open Source Software, 3(23):430:1, March 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00430.

Frost:2016:PGJ

[Fro16] Simon D. W. Frost. Gillespie.jl: Stochastic simulation algorithm in Julia. Journal of Open Source Software, 1(3): 42:1, July 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00042.

Frost:2018:PPJ

[Fro18] Jarvist Moore Frost. PolaronMobility.jl: Implementation of the Feynman variational polaron model. Journal of Open Source Software, 3(28):566:1–566:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00566.

Ferraris:2017:PBM

[FSV^{+}17] Sebastiano Ferraris, Dzhoshkun Ismail Shakir, Johannes Van Der Merwe, Willy Gsell, Jan Deprest, and Tom Vercauteren. Bruker2nifti: Magnetic resonance images converter from

                                        Bruker ParaVision to Nifti

format. Journal of Open Source Software, 2(16):354:1–354:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00354.

Fleming:2018:PAA

[FV18] David P. Fleming and Jake

                                        VanderPlas. approxposterior:

Approximate posterior distributions in Python. Journal of Open Source Software, 3(29):781:1–781:2, September 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00781.

Ganz:2016:PRW

[Gan16] Carl Ganz. rintrojs: A wrapper for the Intro.js library. Journal of Open Source Software, 1(6):63:1, October 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00063.

Gasparini:2018:PCR

[Gas18a] Alessandro Gasparini. comorbidity:

An R package for computing comorbidity scores. Journal of Open Source Software, 3(23):648:1–648:3, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00648.

Gasparini:2018:PRS

[Gas18b] Alessandro Gasparini. rsimsum:

Summarise results from Monte Carlo simulation studies. Journal of Open Source Software, 3(26):739:1–739:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00739.

Goyal:2018:PGP

[GF18] Palash Goyal and Emilio Ferrara. GEM: A Python package for graph embedding methods. Journal of Open Source Software, 3(29):876:1–876:2, September 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00876.

Germonprez:2016:PDS

[GGKG16] Matt Germonprez, Thomas Gurney, Sai Uday Shankar Korlimarla, and Robin Gandhi.

                                        DoSOCS: A system for SPDX 2.0

document creation and storage. Journal of Open Source Software, 1(7):38:1, November 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00038.

Ghoshal:2017:PFF

[GHF^{+}17] Devarshi Ghoshal, Valerie Hendrix, William Fox, Sowmya Balasubhramanian, and Lavanya Ramakrishnan. FRIEDA: Flexible robust intelligent elastic data management framework. Journal of Open Source Software, 2(10):164:1–164:3, February 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00164.

Ginn:2018:PQR

[Gin18] Jasper Ginn. qualtRics: retrieve survey data using the Qualtrics API. Journal of Open Source Software, 3(24):690:1, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00690.

Ganose:2018:PSC

[GJS18] Alex M. Ganose, Adam J. Jackson, and David O. Scanlon. sumo: Command-line tools for plotting and analysis of periodic ab initio calculations. Journal of Open Source Software, 3(28):717:1–717:3, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00717.

Giraud:2016:PCC

[GL16] Timothée Giraud and Nicolas Lambert. cartography: Create and integrate maps in your R workflow. Journal of Open Source Software, 1(4):54:1–54:2, August 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00054.

Goldsmith:2016:PVC

[Gol16] Jeff Goldsmith. vbvs.concurrent:

Fitting methods for the functional linear concurrent model. Journal of Open Source Software, 1(8):141:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00141.

Govan:2018:PBI

[Gov18] Paul B. Govan. BayesianNetwork: Interactive Bayesian network modeling and analysis. Journal of Open Source Software, 3(21):425:1–425:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00425.

Gray:2018:PQP

[Gra18] Johnnie Gray. quimb: A python package for quantum information and manybody calculations. Journal of Open Source Software, 3 (29):819:1–819:3, September 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00819.

Greenhill:2016:PPP

[Gre16] Simon J. Greenhill. Phylogemetric: A Python library for calculating phylogenetic network metrics. Journal of Open Source Software, 1(2):28:1, June 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00028.

Green:2018:PDP

[Gre18] Gregory M. Green. dustmaps: A Python interface for maps of interstellar dust. Journal of Open Source Software, 3(26):695:1–695:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00695.

Grieve:2017:PSE

[Gri17] Stuart W. D. Grieve. spatial-efd:

A spatial-aware implementation of elliptical Fourier analysis. Journal of Open Source Software, 2(11):189:1–189:2, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00189.

Gerosa:2017:PFA

[GV17] Davide Gerosa and Michele Vallisneri. filltex: Automatic queries to ADS and IN-SPIRE databases to fill La-Tex bibliography. Journal of Open Source Software, 2(13): 222:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00222.

Garay:2018:PCA

[GVM18] Pablo G. Garay, Jorge A.

Vila, and Osvaldo A. Martin. CheSweet: An application to predict glycan’s chemicals shifts. Journal of Open Source Software, 3(21):488:1–488:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00488.

Gieseke:2018:PPP

[GWM18] Robert Gieseke, Sven N. Willner, and Matthias Mengel. Pymagicc: A Python wrapper for the simple climate model MAGICC. Journal of Open Source Software, 3(22):516:1– 516:3, February 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00516.

Goldbaum:2018:PUH

[GZT^{+}18] Nathan J. Goldbaum, John A. ZuHone, Matthew J. Turk, Kacper Kowalik, and Anna L. Rosen. unyt: Handle, manipulate, and convert data with units in Python. Journal of Open Source Software, 3(28):809:1–809:11, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00809.

Haghighi:2017:PPP

[Hag17]  Sepand Haghighi.  Pyrgg:

Python Random Graph Generator. Journal of Open Source Software, 2(17):331:1–331:2, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00331.

Haghighi:2018:POO

[HAHR18] Sepand Haghighi, Kasra Askari, Sarmin Hamidi, and Mohammad Mahdi Rahimi. OPEM: Open source PEM cell simulation tool. Journal of Open Source Software, 3(27):676:1– 676:4, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00676.

Halterman:2017:PMF

[Hal17] Andrew Halterman. Mordecai:

Full text geoparsing and event geocoding. Journal of Open Source Software, 2(9):91:1, January 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00091.

Handley:2018:PFP

[Han18] Will Handley. fgivenx: A Python package for functional posterior plotting. Journal of Open Source Software, 3(28):849:1–849:4, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00849.

Harris:2017:EUP

[Har17a] David J. Harris. An easy-to-use p5.js 3D object picker for visual artists. Journal of Open Source Software, 2(20):475:1–475:2, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00475.

Harrison:2017:PVA

[Har17b] Paul Francis Harrison. Varistran:

Anscombe’s variance stabilizing transformation for RNAseq gene expression data. Journal of Open Source Software, 2 (16):257:1, August 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00257.

Holtgrewe:2016:PVP

[HB16] Manuel Holtgrewe and Dieter Beule. VCFPy: a Python 3 library with good support for both reading and writing VCF. Journal of Open Source Software, 1(6):85:1, October 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00085.

Hejazi:2017:PBT

[HCH17] Nima S. Hejazi, Weixin Cai, and Alan E. Hubbard. biotmle: Targeted learning for biomarker discovery. Journal of Open Source Software, 2(15):295:1– 295:4, July 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00295.

Hiebert:2018:PCC

[HCM^{+}18] James Hiebert, Alex J. Cannon, Trevor Murdock, Stephen Sobie, and Arelia Werner. ClimDown: Climate downscaling in R. Journal of Open Source Software, 3(22):360:1, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00360.

Hobson:2017:PDR

[HDL17] Tanner C. Hobson, Mathieu Doucet, and Ricardo M. Ferraz Leal. Django remote submission. Journal of Open Source Software, 2(16):366:1–366:2, August 2017. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00366.

Hermes:2017:HBC

[Her17] Danny Hermes. Helper for bézier curves, triangles, and higher order objects. Journal of Open Source Software, 2(16):267:1–267:3, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00267.

Hernandez:2016:PXP

[HF16] Guillermo Hernández and Francisco Fernández. xpecgen: A program to calculate Xray spectra generated in tungsten anodes. Journal of Open Source Software, 1(7):62:1– 62:2, November 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00062.

Heusser:2017:PQP

[HFF^{+}17] Andrew C. Heusser, Paxton C.

Fitzpatrick, Campbell E. Field, Kirsten Ziman, and Jeremy R. Manning. Quail: A Python toolbox for analyzing and plotting free recall data. Journal of Open Source Software, 2(18):424:1–424:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00424.

Harpole:2016:PRR

[HH16] Alice Harpole and Ian Hawke. R3D2: Relativistic Reactive Riemann problem solver for Deflagrations and Detonations. Journal of Open Source Software, 1(1):16:1, May 2016. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00016.

Holmgren:2018:PPP

[HHM18] William F. Holmgren, Clifford W. Hansen, and Mark A.

                                        Mikofski. pvlib python: a

python package for modeling solar energy systems. Journal of Open Source Software, 3(29):884:1–884:3, September 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00884.

Hernandez:2017:PMD

[HHSP17] Carlos X. Hernández, Matthew P. Harrigan, Mohammad M. Sultan, and Vijay S. Pande. MSMExplorer: Data visualizations for biomolecular dynamics. Journal of Open Source Software, 2(12):188:1–188:2, April 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00188.

Higgins:2017:PVF

[HHV17] Joshua Higgins, Violeta Holmes, and Colin Venters. VCC: A framework for building containerized reproducible cluster software environments. Journal of Open Source Software, 2 (11):208:1, March 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00208.

Hickey:2016:RMG

[Hic16] Peter F. Hickey. Representation and manipulation of genomic tuples in R. Journal of Open Source Software, 1(1): 20:1, May 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00020.

Hinton:2016:PC

[Hin16] Samuel Hinton. ChainConsumer.

Journal of Open Source Software, 1(4):45:1–45:2, August 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00045.

Hirsch:2018:PPD

[Hir18] Michael Hirsch. PyMap3D: 3-D coordinate conversions for terrestrial and geospace environments. Journal of Open Source Software, 3(23):580:1–580:2, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00580.

Hudak:2018:OOW

[HJC^{+}18] Dave Hudak, Doug Johnson, Alan Chalker, Jeremy Nicklas, Eric Franz, Trey Dockendorf, and Brian L. McMichael. Open OnDemand: A web-based client portal for HPC centers. Journal of Open Source Software, 3(25):622:1–622:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00622.

Haghighi:2018:PPM

[HJHZ18] Sepand Haghighi, Masoomeh Jasemi, Shaahin Hessabi, and Alireza Zolanvari. PyCM: Multiclass confusion matrix library in Python. Journal of Open Source Software, 3(25):729:1– 729:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00729.

Hazelton:2017:PPI

[HJPB17] Bryna J. Hazelton, Daniel C. Jacobs, Jonathan C. Pober, and Adam P. Beardsley. pyuvdata: an interface for astronomical interferometeric datasets in python. Journal of Open Source Software, 2(10):140:1, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00140.

Hoffimann:2018:PGJ

[Hof18] Júlio Hoffimann. GeoStats.jl — high-performance geostatistics in Julia. Journal of Open Source Software, 3(24): 692:1–692:4, April 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.

org/papers/10.21105/joss.
00692.

Honorato:2016:PCP

[Hon16] Rodrigo V. Honorato. CAZy-parser a way to extract information from the Carbohydrate-Active enZYmes database. Journal of Open Source Software, 1(8):53:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00053.

Honkonen:2017:HDI

[Hon17] Ilja Honkonen. High-dimensional integrator. Journal of Open Source Software, 2(20):437:1, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00437.

Hope:2017:PIM

[Hop17] Gaute Hope. IBCAO py: A matplotlib library for using the International Bathymetric Chart of the Arctic Ocean with cartopy and matplotlib. Journal of Open Source Software, 2(13):250:1–250:2, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00250.

Howard:2017:VBA

[How17] James P. Howard II. Virtual bumblebees artificial life simulation. Journal of Open Source Software, 2(13):256:1, May 2017. CODEN ????

ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00256.

Howard:2018:PAR

[How18] James P. Howard II. Phonetic algorithms in R. Journal of Open Source Software, 3(22):480:1–480:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00480.

Hernandez:2017:PMI

[HP17] Carlos X. Hernández and Vijay S. Pande. MDEntropy: Information-theoretic analyses for molecular dynamics. Journal of Open Source Software, 2(19):427:1–427:2, November 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00427.

Hagan:2018:PPO

[HTT18] David H. Hagan, Andrew Tolmie, and Jakub Trochim. py-opc: operate the Alphasense OPC-N2 from a raspberry pi or other popular microcontrollers/microcomputers. Journal of Open Source Software, 3(26):782:1–782:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00782.

Herman:2017:PSO

[HU17] Jon Herman and Will Usher. SALib: An open-source Python library for sensitivity analysis. Journal of Open Source Software, 2(9):97:1–97:2, January 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00097.

Hucka:2018:PNN

[Huc18a] Michael Hucka. Nostril: A nonsense string evaluator written in Python. Journal of Open Source Software, 3(25):596:1– 596:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00596.

Hucka:2018:PSS

[Huc18b] Michael Hucka. Spiral: splitters for identifiers in source code files. Journal of Open Source Software, 3(24):653:1– 653:3, April 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00653.

Hughes:2016:PPR

[Hug16] Sean M. Hughes. plater:

Read, tidy, and display data from microtiter plates. Journal of Open Source Software, 1(7):106:1, November 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00106.

Hughes:2018:PMP

[Hug18] Momar G-O Hughes. MCycle:

A Python package for 1D sizing and analysis of thermodynamic power cycles. Journal of Open Source Software, 3(28):710:1–710:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00710.

Ibarra-Espinosa:2018:PER

[IESdF18] Sergio Ibarra-Espinosa, Daniel Schuch, and Edmilson Dias de Freitas. eixport: An R package to export emissions to atmospheric models. Journal of Open Source Software, 3(24):607:1–607:4, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00607.

Innes:2018:PFE

[Inn18] Mike Innes. Flux: Elegant machine learning with Julia. Journal of Open Source Software, 3(25):602:1, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00602.

Iravanian:2018:PFT

[Ira18] Shahriar Iravanian. fib-tf: A TensorFlow-based cardiac electrophysiology simulator. Journal of Open Source Software, 3(26):719:1–719:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00719.

Izzo:2018:VMT

[ISMA18] Richard Izzo, David Steinman, Simone Manini, and Luca Antiga. The vascular modeling toolkit: A Python library for the analysis of tubular structures in medical images. Journal of Open Source Software, 3(25):745:1–745:5, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00745.

Jacobs:2016:PGR

[JA16] Christian T. Jacobs and Alexandros Avdis. Git-RDM: A research data management plugin for the Git version control system. Journal of Open Source Software, 1(2):29:1–29:2, June 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00029.

Jensen:2017:POA

[JCS17] Kristian Jensen, Joao G. R.

Cardoso, and Nikolaus Sonnenschein. Optlang: An algebraic modeling language for mathematical optimization. Journal of Open Source Software, 2(9):139:1, January 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00139.

James:2018:PDP

[JEC18] Ryan G. James, Christopher J.

Ellison, and James P. Crutchfield. dit: a Python package for discrete information theory. Journal of Open Source Software, 3(25):738:1–738:3, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00738.

Jimenez:2017:PPB

[JG17] José Jiménez and Josep Ginebra. pyGPGO: Bayesian optimization for Python. Journal of Open Source Software, 2(19):431:1–431:3, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00431.

Jackson:2018:PGB

[JGR^{+}18] Adam J. Jackson, Alex M. Ganose, Anna Regoutz, Russell G. Egdell, and David O. Scanlon. Galore: Broadening and weighting for simulation of photoelectron spectroscopy. Journal of Open Source Software, 3(26):773:1– 773:6, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00773.

Jones:2016:PEE

[JL16] Zachary M. Jones and Fridolin J. Linder. edarf: Exploratory Data Analysis using Random Forests. Journal of Open Source Software, 1(6):92:1– 92:4, October 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.

org/papers/10.21105/joss.
00092.

Jones:2018:PLL

[JNM18] Andrew T. Jones, Hien D.

Nguyen, and Geoffrey J. McLachlan. logKDE: logtransformed kernel density estimation. Journal of Open Source Software, 3(28):870:1– 870:3, August 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00870.

Joyce:2017:PLI

[Joy17] P. James Joyce. Lcopt — an interactive tool for creating fully parameterised life cycle assessment (LCA) foreground models. Journal of Open Source Software, 2(16):339:1, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00339.

Kortschak:2017:PBN

[KA17] R. Daniel Kortschak and David L. Adelson. bíogo/ncbi: interfaces to NCBI services for the Go language. Journal of Open Source Software, 2(18):234:1–234:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00234.

Kane:2017:PBR

[Kan17] Michael Kane. bittrex: An R client for the Bittrex Crypto-Currency Exchange. Journal of Open Source Software, 2(17):402:1, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00402.

Karpel:2018:PIG

[Kar18] Joshua T. Karpel. IDESolver: a general purpose integrodifferential equation solver. Journal of Open Source Software, 3(21):542:1–542:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00542.

Krebber:2018:PMP

[KB18] Manuel Krebber and Henrik Barthels. MatchPy: Pattern matching in Python. Journal of Open Source Software, 3(26):670:1–670:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00670.

Kearney:2017:PEM

[Kea17] Kelly A. Kearney. ecopath matlab: A Matlab-based implementation of the Ecopath food web algorithm. Journal of Open Source Software, 2(9): 64:1–64:2, January 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00064.

Klebel:2018:PJI

[Kle18] Thomas Klebel. jstor: Import and analyse data from scientific texts. Journal of Open Source Software, 3(28):883:1–883:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00883.

Kessler:2017:PEL

[KM17a] Travis Kessler and John Hunter Mack. ECNet: Large scale machine learning projects for fuel property prediction. Journal of Open Source Software, 2(17):401:1–401:2, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00401.

Kleinberg:2017:WBT

[KM17b] Bennett Kleinberg and Maximilian Mozes. Web-based text anonymization with Node.js: Introducing NETANOS (Named entity-based Text Anonymization for Open Science). Journal of Open Source Software, 2(14):293:1–293:2, June 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00293.

Krishnan:2017:PCG

[KMB17] Anush Krishnan, Olivier Mesnard, and Lorena A. Barba. cuIBM: a GPU-based immersed boundary method code. Journal of Open Source Software, 2(15):301:1, July 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00301.

Kironde:2017:PRD

[KMG^{+}17] Henry Kironde, Benjamin D. Morris, Akash Goel, Andrew Zhang, Akshay Narasimha, Shivam Negi, David J. Harris, Deborah Gertrude Digges, Kapil Kumar, Amritanshu Jain, Kunal Pal, Kevinkumar Amipara, Prabh Simran Singh Baweja, and Ethan P. White. Retriever: Data retrieval tool. Journal of Open Source Software, 2(19):451:1, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00451.

Kortschak:2018:PAG

[Kor18] R. Daniel Kortschak. arrgh: a Go interface to the OpenCPU R server system. Journal of Open Source Software, 3(21): 517:1, January 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00517.

Kortschak:2017:PBH

[KPA17] R. Daniel Kortschak, Brent S. Pedersen, and David L. Adelson. bíogo/hts: high throughput sequence handling for the Go language. Journal of Open Source Software, 2(10):168:1, February 2017. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00168.

Kamppi:2017:PKG

[KPV^{+}17] Antti Kamppi, Esko Pekkarinen, Janne Virtanen, Joni-Matti Määttä, Juho Järvinen, Lauri Matilainen, Mikko Teuho, and Timo D. Hämäläinen. Kactus2: A graphical EDA tool built on the IP-XACT standard. Journal of Open Source Software, 2(13):151:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00151.

Kortschak:2017:PBS

[KSMA17] R. Daniel Kortschak, Josh Bleecher Snyder, Manolis Maragkakis, and David L. Adelson. bíogo: a simple high-performance bioinformatics toolkit for the Go language. Journal of Open Source Software, 2(10):167:1, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00167.

LaZerte:2018:PWD

[LA18] Stefanie E. LaZerte and Sam Albers. weathercan: Download and format weather data from Environment and Climate Change Canada. Journal of Open Source Software, 3(22):571:1–571:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00571.

Landau:2018:PDR

[Lan18] William Michael Landau. The drake R package: a pipeline toolkit for reproducibility and high-performance computing. Journal of Open Source Software, 3(21):550:1, January 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00550.

Laros:2018:GBF

[Lar18] Jeroen F. J. Laros. General binary file parser. Journal of Open Source Software, 3(26): 766:1, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00766.

Latham:2017:PAM

[Lat17] Shane J. Latham. array split: Multi-dimensional array partitioning. Journal of Open Source Software, 2(17):373:1– 373:2, September 2017. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00373.

Lauwens:2017:PRC

[Lau17] Ben Lauwens. ResumableFunctions: C# sharp style generators for Julia. Journal of Open Source Software, 2(18):400:1–400:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00400.

Laurinec:2018:PTR

[Lau18] Peter Laurinec. TSrepr R package: Time series representations. Journal of Open Source Software, 3(23):577:1–577:2, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00577.

Lockley:2017:PXE

[LBČ17] Steve Lockley, Claudio Benghi, and Martin Černý. Xbim.Essentials: a library for interoperable building information applications. Journal of Open Source Software, 2(20):473:1–473:3, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00473.

Lang:2017:PBT

[LBS17] Michel Lang, Bernd Bischl, and Dirk Surmann. batchtools: Tools for R to work on batch systems. Journal of Open Source Software, 2(10):135:1– 135:2, February 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00135.

Louppe:2016:PCL

[LCP16] Gilles Louppe, Kyle Cranmer, and Juan Pavez. carl: a likelihood-free inference toolbox. Journal of Open Source Software, 1(1):11:1, May 2016. CODEN ???? ISSN 2475-9066.

URL  http://joss.theoj.

                                        org/papers/10.21105/joss.
                                        00011.

LeBeau:2018:PHC

[LeB18a] Brandon LeBeau. highlightHTML: CSS formatting of R Markdown documents. Journal of Open Source Software, 3(21):185:1– 185:2, January 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00185.

LeBeau:2018:PPS

[LeB18b] Brandon LeBeau. pdfsearch: Search tools for PDF files. Journal of Open Source Software, 3(27):668:1–668:2, July 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00668.

Lee:2016:PFF

[Lee16] Seward Lee. finreportr: Financial data from U.S. Securities and Exchange Commission. Journal of Open Source Software, 1(8):119:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00119.

Levy:2016:PGI

[Lev16] Michael A. Levy. gwdegree: Improving interpretation of geometrically-weighted degree estimates in exponential random graph models. Journal of Open Source Software, 1(3): 36:1, July 2016. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00036.

Lee:2018:PDD

[LGH^{+}18] Jeongseok Lee, Michael X.

Grey, Sehoon Ha, Tobias Kunz, Sumit Jain, Yuting Ye, Siddhartha S. Srinivasa, Mike Stilman, and C. Karen Liu. DART: Dynamic animation and robotics toolkit. Journal of Open Source Software, 3(22):500:1–500:3, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00500.

Lindsay:2018:PMF

[LH18] Alexander Lindsay and Kathryn Huff. Moltres: finite element based simulation of molten salt reactors. Journal of Open Source Software, 3(21):298:1–298:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00298.

Light:2017:PMS

[Lig17] Roger A. Light. Mosquitto:

server and client implementation of the MQTT protocol. Journal of Open Source Software, 2(13):265:1–265:2, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00265.

Lin:2018:PMP

[LIK18] Jiao Y. Y. Lin, Fahima Islam, and Max Kresh. Multiphonon: Phonon density of states tools for inelastic neutron scattering powder data. Journal of Open Source Software, 3(21):440:1–440:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00440.

Ludecke:2018:PGT

[Lüd18a] Daniel Lüdecke. ggeffects: Tidy data frames of marginal effects from regression models. Journal of Open Source Software, 3(26):772:1–772:5, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00772.

Ludecke:2018:PSD

[Lüd18b] Daniel Lüdecke. sjmisc: Data and variable transformation functions. Journal of Open Source Software, 3(26):754:1– 754:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00754.

Madan:2016:PPM

[Mad16] Christopher R. Madan. Prism: Multiple spline regression with regularization, dimensionality reduction, and feature selection. Journal of Open Source Software, 1(3):31:1–31:3, July 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00031.

Makela:2016:PIL

[Mäk16] Eetu Mäkelä. LAS: an integrated language analysis tool for multiple languages. Journal of Open Source Software, 1(6):35:1–35:2, October 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00035.

Makowski:2018:PPP

[Mak18] Dominique Makowski. The psycho package: an efficient and publishing-oriented workflow for psychological science. Journal of Open Source Software, 3(22):470:1–470:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00470.

Margolis:2017:PSP

[Mar17] Benjamin W. L. Margolis.

SimuPy: A Python framework for modeling and simulating dynamical systems. Journal of Open Source Software, 2(17):396:1, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00396.

Mayer:2017:PNP

[May17] Andreas Mayer. Noisyopt: A Python library for optimizing noisy functions. Journal of Open Source Software, 2(13): 258:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00258.

Mahar:2017:PHP

[MB17a] Sara Mahar and Matthew Bellis. hmis: A python tool to visualize and analyze HMIS data. Journal of Open Source Software, 2(18):384:1–384:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00384.

Murray:2017:PLC

[MB17b] Kevin D. Murray and Justin O. Borevitz. libqcpp: A C++14 sequence quality control library. Journal of Open Source Software, 2(13):232:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00232.

Mullen:2018:PUH

[MB18] Lincoln A. Mullen and Jordan Bratt. USAboundaries: Historical and contemporary boundaries of the United States of America. Journal of Open Source Software, 3(23):314:1– 314:2, March 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00314.

Mullen:2018:FCT

[MBK^{+}18] Lincoln A. Mullen, Kenneth Benoit, Os Keyes, Dmitry Selivanov, and Jeffrey Arnold. Fast, consistent tokenization of natural language text. Journal of Open Source Software, 3(23):655:1–655:3, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00655.

Molnar:2018:PIR

[MCB18] Christoph Molnar, Giuseppe Casalicchio, and Bernd Bischl. iml: An R package for interpretable machine learning. Journal of Open Source Software, 3(26):786:1–786:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00786.

McFee:2016:PRE

[McF16] Brian McFee. resampy: efficient sample rate conversion in Python. Journal of Open Source Software, 1(8):125:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00125.

McLean:2017:PRI

[McL17] Mathew W. McLean. RefManageR:

Import and manage BibTeX and BibLaTeX references in R. Journal of Open Source Software, 2(16):338:1–338:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00338.

Makowski:2017:PNP

[MD17] Dominique Makowski and Léo

                                        Dutriaux. Neuropsydia.py: A

Python module for creating experiments, tasks and questionnaires. Journal of Open Source Software, 2(19):259:1–259:2, November 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00259.

Morris:2018:PAA

[MDW18] Brett M. Morris and Trevor Dorn-Wallenstein. aesop: ARC Echelle Spectroscopic Observation Pipeline. Journal of Open Source Software, 3 (28):854:1, August 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00854.

Marzolin:2018:FGG

[MG18a] Gilbert Marzolin and Olivier Gimenez. Fitting a Gamma– Gompertz survival model to capture-recapture data collected on free-ranging animal populations. Journal of Open Source Software, 3(21):216:1– 216:3, January 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00216.

Mikhael:2018:PMM

[MG18b] Shadia Mikhael and Calum

                                        Gray. Masks2Metrics (M2M):

A Matlab toolbox for gold standard morphometrics. Journal of Open Source Software, 3(22):436:1–436:3, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00436.

Moore:2018:PSS

[MH18] Jason K. Moore and Mont

Hubbard.  skijumpdesign:

A ski jump design tool for specified equivalent fall height. Journal of Open Source Software, 3(28):818:1–818:3, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00818.

McInnes:2017:PHH

[MHA17] Leland McInnes, John Healy, and Steve Astels. hdbscan: Hierarchical density based clustering. Journal of Open Source Software, 2(11):205:1–205:2, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00205.

McGibbon:2016:POH

[MHH^{+}16] Robert T. McGibbon, Carlos X. Hernández, Matthew P. Harrigan, Steven Kearnes, Mohammad M. Sultan, Stanisław Jastrzebski, Brooke E. Husic, and Vijay S. Pande. Osprey: Hyperparameter optimization for machine learning. Journal of Open Source Software, 1(5):34:1–34:2, September 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00034.

McInnes:2018:PUU

[MHSG18] Leland McInnes, John Healy, Nathaniel Saul, and Lukas Großberger. UMAP: Uniform manifold approximation and projection. Journal of Open Source Software, 3(29):861:1– 861:2, September 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00861.

Miranda:2018:PPR

[Mir18] Lester James V. Miranda. PySwarms: a research toolkit for particle swarm optimization in Python. Journal of Open Source Software, 3(21):433:1– 433:2, January 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00433.

Malengier:2018:POH

[MKT^{+}18] Benny Malengier, Pavol Kišon, James Tocknell, Claas Abert, Florian Bruckner, and Marc-Antonio Bisotti. ODES: a high level interface to ODE and DAE solvers. Journal of Open Source Software, 3(22):165:1–165:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00165.

Madnani:2016:PRC

[ML16] Nitin Madnani and Anastassia Loukina. RSMTool: collection of tools building and evaluating automated scoring models. Journal of Open Source Software, 1(3):33:1, July 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00033.

Moerman:2018:PGG

[Moe18] Kevin M. Moerman. GIBBON:

The Geometry and Image-Based Bioengineering add-On. Journal of Open Source Software, 3(22):506:1–506:4, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00506.

Morgan:2017:PBB

[Mor17a] Benjamin J. Morgan. bsym:

A basic symmetry module. Journal of Open Source Software, 2(16):370:1–370:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00370.

Morgan:2017:PLP

[Mor17b] Benjamin J. Morgan. lattice mc:

A Python lattice-gas Monte Carlo module. Journal of Open Source Software, 2(13):247:1– 247:2, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00247.

Moritz:2017:TDS

[Mor17c] Dominik Moritz. Text detection in screen images with a convolutional neural network. Journal of Open Source Software, 2(15):235:1, July 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00235.

Morel:2018:PGG

[Mor18] Pierre Morel. Gramm: grammar of graphics plotting in Matlab. Journal of Open Source Software, 3(23):568:1–568:4, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00568.

McCormick:2017:PSS

[MPFRA17] Matthew M. McCormick, Mark L. Palmeri, Jean-Christophe Fillion-Robin, and Stephen Aylward.

SlicerITKUltrasound: A 3D

slicer extension for scan conversion of B-mode and nextgeneration ultrasound imaging modalities. Journal of Open Source Software, 2(10):153:1, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00153.

Mogensen:2018:POM

[MR18a] Patrick K. Mogensen and Asbjørn N. Riseth. Optim: A mathematical optimization package for Julia. Journal of Open Source Software, 3(24):615:1–615:3, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00615.

Mulryne:2018:PPP

[MR18b] David J. Mulryne and John W.

Ronayne. PyTransport: A

Python package for the calculation of inflationary correlation functions. Journal of Open Source Software, 3(23):494:1–494:2, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00494.

Michaels:2016:PHR

[MS16] Jonathan A. Michaels and Hansjörg Scherberger. hebbRNN: A reward-modulated Hebbian learning rule for recurrent neural networks. Journal of Open Source Software, 1(5):60:1– 60:2, September 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00060.

McGinnis:2018:CEP

[MSSH18] William D. McGinnis, Chapman Siu, Andre S., and Hanyu Huang. Category encoders: a

scikit-learn-contrib pack-

age of transformers for encoding categorical data. Journal of Open Source Software, 3(21):501:1–501:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00501.

Murray:2018:PPP

[Mur18] Steven G. Murray. powerbox: A Python package for creating structured fields with isotropic power spectra. Journal of Open Source Software, 3(28):850:1–850:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00850.

Mutel:2017:PBO

[Mut17a] Chris Mutel. Brightway: An open source framework for life cycle assessment. Journal of Open Source Software, 2(12):236:1–236:2, April 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00236.

Mutel:2017:PPG

[Mut17b] Chris Mutel. Pandarus: GIS toolkit for regionalized life cycle assessment. Journal of Open Source Software, 2(13): 244:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00244.

Moore:2018:POS

[MvdB18] Jason K. Moore and Antonie van den Bogert. opty: Software for trajectory optimization and parameter identification using direct collocation. Journal of Open Source Software, 3(21):300:1–300:4, January 2018. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00300.

Mirzaev:2018:PEP

[MWS18] Inom Mirzaev, Drew F. K.

Williamson, and Jacob G. Scott. egtplot: A Python package for three-strategy evolutionary games. Journal of Open Source Software, 3(26): 735:1–735:4, June 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00735.

Mount:2018:PVR

[MZ18] John Mount and Nina Zumel.

The vtreat R package: a statistically sound data processor for predictive modeling. Journal of Open Source Software, 3(23):584:1–584:2, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00584.

Nunez-Elizalde:2018:PCS

[NEGZG18] Anwar O. Nunez-Elizalde, James S. Gao, Tianjiao Zhang, and Jack L. Gallant. cottoncandy: scientific python package for easy cloud storage. Journal of Open Source Software, 3(28):890:1–890:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00890.

Naecker:2017:PPP

[NMGB17] Benjamin Naecker, Niru Maheswaranathan, Surya Ganguli, and Stephen Baccus. Pyret: A Python package for analysis of neurophysiology data. Journal of Open Source Software, 2(9):137:1, January 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00137.

Nolan:2017:PFR

[NPP17] Rory Nolan and Sergi Padilla-

                                        Parra. filesstrings: An R

package for file and string manipulation. Journal of Open Source Software, 2(14):260:1– 260:2, June 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00260.

Nolan:2018:PIR

[NPP18] Rory Nolan and Sergi Padilla-Parra. ijtiff: An R package providing TIFF I/O for ImageJ users. Journal of Open Source Software, 3(23):633:1–633:2, March 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00633.

Neches:2018:PSF

[NS18] Russell Y. Neches and Camille Scott. SuchTree: Fast, threadsafe computations with phylogenetic trees. Journal of Open Source Software, 3(27):678:1– 678:3, July 2018. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00678.

Norris:2017:PPA

[NSB17] Clayton Norris, Philip Schrodt, and John Beieler. PETRARCH2: Another event coding program. Journal of Open Source Software, 2(9):133:1, January 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00133.

Niskanen:2017:PAT

[NSL17] Julia Niskanen, Elina Salmela, and Hannes Lohi. AncesTrim — a tool for trimming complex pedigrees. Journal of Open Source Software, 2(11):179:1, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00179.

Okada:2017:FGI

[OCT^{+}17] Ai Okada, Kenichi Chiba, Hiroko Tanaka, Satoru Miyano, and Yuichi Shiraishi. A framework for generating interactive reports for cancer genome analysis. Journal of Open Source Software, 2(20):457:1–457:4, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00457.

Odell:2018:PNA

[Ode18] Evan Odell. nomisr: Access Nomis UK labour market data. Journal of Open Source Software, 3(27):859:1– 859:2, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00859.

Owan:2018:PCO

[ODP18] Parker Owan, Cameron Devine, and W. Tony Piaskowy. CoreRobotics: An object-oriented C++ library with cross-language wrappers for cross-platform robot control. Journal of Open Source Software, 3(22):489:1– 489:2, February 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00489.

Ortiz-Gil:2018:MFC

[OGBC18] Amelia Ortiz-Gil and Jordi Burguet-Castell. Mapelia and friends: create 3D models from maps. Journal of Open Source Software, 3(25):660:1– 660:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00660.

Oliver:2018:PCW

[OSS18] Hilary J. Oliver, Matthew Shin, and Oliver Sanders. Cylc: A workflow engine for cycling systems. Journal of Open Source Software, 3(27):737:1– 737:2, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00737.

Ostrouchov:2018:PPA

[OZW18] Christopher Ostrouchov, Yanwen Zhang, and William J. Weber. pysrim: Automation, analysis, and plotting of SRIM calculations. Journal of Open Source Software, 3(28):829:1–829:3, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00829.

Pastell:2017:PWJ

[Pas17]  Matti Pastell.  Weave.jl:

Scientific reports using Julia. Journal of Open Source Software, 2(11):204:1, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00204.

Pearse:2018:PSD

[PC18] William D. Pearse and Scott A.

Chamberlain. Suppdata: Downloading supplementary data from published manuscripts. Journal of Open Source Software, 3(25):721:1–721:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00721.

Padgham:2017:PB

[PE17] Mark Padgham and Richard Ellison. bikedata. Journal of Open Source Software, 2(20):471:1, December 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00471.

Page:2017:PSF

[PHSK17] Andrew J. Page, Martin Hunt, Torsten Seemann, and Jacqueline A. Keane. SaffronTree: Fast, reference-free pseudophylogenomic trees from reads or contigs. Journal of Open Source Software, 2(13):243:1– 243:2, May 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00243.

Pitkin:2018:PPP

[Pit18] Matthew Pitkin. psrqpy: a python interface for querying the ATNF pulsar catalogue. Journal of Open Source Software, 3(22):538:1–538:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00538.

Ponge:2016:EG

[PLL^{+}16] Julien Ponge, Yannick Loiseau, Frédéric Le Mouël, Nicolas Stouls, Philippe Charrière, Daniel Petisme, Sylvain Desgrais, and Franck Verrot. Eclipse Golo. Journal of Open Source Software, 1(8):93:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00093.

Padgham:2017:PO

[PLSR17] Mark Padgham, Robin Lovelace, Maëlle Salmon, and Bob Rudis. osmdata. Journal of Open Source Software, 2(14):305:1, June 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00305.

Porter:2018:PJE

[Por18] Michael T. Porter. js-emass:

A flexible JavaScript implementation of the emass algorithm. Journal of Open Source Software, 3(28):869:1, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00869.

Pfenninger:2018:PCM

[PP18] Stefan Pfenninger and Bryn Pickering. Calliope: a multiscale energy systems modelling framework. Journal of Open Source Software, 3 (29):825:1–825:2, September 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00825.

Polimis:2017:CIR

[PRH17] Kivan Polimis, Ariel Rokem, and Bryna Hazelton. Confidence intervals for random forests in Python. Journal of Open Source Software, 2(19):124:1–124:4, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00124.

Page:2016:PGP

[PST^{+}16] Andrew J. Page, Sascha Steinbiss, Ben Taylor, Torsten Seemann, and Jacqueline A. Keane. GFF3toEMBL: Preparing annotated assemblies for submission to EMBL. Journal of Open Source Software, 1(6):80:1, October 2016. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00080.

Page:2016:MST

[PTK16] Andrew J. Page, Ben Taylor, and Jacqueline A. Keane. Multilocus sequence typing by blast from de novo assemblies against PubMLST. Journal of Open Source Software, 1(8):118:1–118:2, December 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00118.

Price-Whelan:2017:PGP

[PW17] Adrian M. Price-Whelan. Gala: A Python package for galactic dynamics. Journal of Open Source Software, 2(18):388:1– 388:2, October 2017. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00388.

Price-Whelan:2017:PSU

[PWFM17] Adrian M. Price-Whelan and Daniel Foreman-Mackey. schwimmbad: A uniform interface to parallel processing pools in Python.

Journal of Open Source Software, 2(17):357:1–357:2, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00357.

Raab:2016:PEU

[Raa16] Markus Raab. Elektra: universal framework to access configuration parameters. Journal of Open Source Software, 1(8):44:1–44:2, December 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00044.

Raschka:2017:PBW

[Ras17] Sebastian Raschka. BioPandas:

Working with molecular structures in pandas DataFrames. Journal of Open Source Software, 2(14):279:1, June 2017. CODEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00279.

Raschka:2018:PMP

[Ras18] Sebastian Raschka. MLxtend:

Providing machine learning and data science utilities and extensions to Python’s scientific computing stack. Journal of Open Source Software, 3(24):638:1–638:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00638.

Roubeyrie:2018:PWP

[RC18] Lionel Roubeyrie and Sébastien Celles. Windrose: A Python

                                        Matplotlib, Numpy library to

manage wind and pollution data, draw windrose. Journal of Open Source Software, 3(29):268:1–268:5, September 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00268.

Rampin:2016:PRR

[RCS^{+}16] Rémi Rampin, Fernando Chirigati, Dennis Shasha, Juliana Freire, and Vicky Steeves. ReproZip: The reproducibility packer. Journal of Open Source Software, 1(8):107:1–107:3, December 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00107.

Reinecke:2018:PGF

[Rei18] Robert Reinecke. G^{3}M-f a global gradient-based groundwater modelling framwork. Journal of Open Source Software, 3(22):548:1–548:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00548.

Renard:2017:PHO

[Ren17]  Philippe Renard.  Hytool:

an open source MATLAB toolbox for the interpretation of hydraulic tests using analytical solutions. Journal of Open Source Software, 2(19):441:1–441:3, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00441.

Rognes:2017:PCA

[RFF^{+}17] Marie E. Rognes, Patrick E.

Farrell, Simon W. Funke, Johan E. Hake, and Molly M. C. Maleckar. cbcbeat: an adjointenabled framework for computational cardiac electrophysiology. Journal of Open Source Software, 2(13):224:1, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00224.

Ramasubramani:2018:PRP

[RG18] Vyas Ramasubramani and Sharon C. Glotzer. rowan: A Python package for working with quaternions. Journal of Open Source Software, 3(27):787:1–787:3, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00787.

Rodenhauser:2018:PPP

[RGZ^{+}18] Anton Rodenhauser, Wilson W. Good, Brian Zenger, Jess Tate, Kedar Aras, Brett Burton, and Rob S. MacLeod. PFEIFER: Preprocessing framework for electrograms intermittently fiducialized from experimental recordings. Journal of Open Source Software, 3(21):472:1–472:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00472.

Reghenzani:2018:PCM

[RMF18] Federico Reghenzani, Giuseppe Massari, and William Fornaciari. chronovise: Measurementbased probabilistic timing analysis framework. Journal of Open Source Software, 3(28): 711:1–711:3, August 2018. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00711.

Romano:2017:PY

[Rom17]  Luca Romano.  YoungTab.

Journal of Open Source Software, 2(13):221:1, May 2017. CODEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00221.

Rose:2018:PCP

[Ros18] Brian E. J. Rose. CLIMLAB: a Python toolkit for interactive, process-oriented climate modeling. Journal of Open Source Software, 3(24):659:1–659:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00659.

Rovigatti:2017:PFE

[Rov17]  Gabriele Rovigatti.  Pro-

duction function estimation in R: The prodest package. Journal of Open Source Software, 2(18):371:1, October 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00371.

Roy:2018:PBS

[RRD^{+}18] Pamphile T. Roy, Sophie Ricci, Romain Dupuis, Robin Campet, Jean-Christophe Jouhaud, and Cyril Fournier. BATMAN: Statistical analysis for expensive computer codes made easy. Journal of Open Source Software, 3(21):493:1–493:2, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00493.

Raamana:2017:HWN

[RS17a] Pradeep Reddy Raamana and Stephen C. Strother. Histogram-weighted networks for feature extraction, connectivity and advanced analysis in neuroscience. Journal of Open Source Software, 2(19):380:1–380:3, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00380.

Raamana:2017:PPC

[RS17b] Pradeep Reddy Raamana and Stephen C. Strother. Python class defining a machine learning dataset ensuring key-based correspondence and maintaining integrity. Journal of Open Source Software, 2(17):382:1, September 2017. CODEN ????

ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00382.

Sloan:2016:PGF

[SAB^{+}16] Zachary Sloan, Danny Arends, Karl W. Broman, Arthur Centeno, Nicholas Furlotte, Harm Nijveen, Lei Yan, Xiang Zhou, Robert W. Williams, and Pjotr Prins. GeneNetwork: framework for web-based genetics. Journal of Open Source Software, 1(2):25:1–25:3, June 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00025.

Standage:2017:PKR

[SAC^{+}17] Daniel Standage, Ali Aliyari, Lisa J. Cohen, Michael R. Crusoe, Tim Head, Luiz Irber, Shannon E. K. Joslin, N. B. Kingsley, Kevin D. Murray, Russell Neches, Camille Scott, Ryan Shean, Sascha Steinbiss, Cait Sydney, and C. Titus Brown. khmer release v2.1: software for biological sequence analysis. Journal of Open Source Software, 2(15):272:1– 272:2, July 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00272.

Strozzi:2017:PPU

[SB17] Francesco Strozzi and Raoul Jean Pierre Bonnal. Pipengine: an ultra light YAML-based pipeline execution engine. Journal of Open Source Software, 2(16):341:1–341:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00341.

Schmeisser:2017:PER

[SBL^{+}17] Andre Schmeißer, Daniel Burkhart, Dominik Linn, Johannes Schnebele, Manuel Ettmüller, Simone Gramsch, and Walter Arne.

Ensight4Matlab: read, pro-

cess, and write files in En-Sight(R) gold format from C++ or MATLAB(R). Journal of Open Source Software, 2(20):217:1, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00217.

Sanderson:2016:PAT

[SC16] Conrad Sanderson and Ryan

Curtin.  Armadillo:  a

template-based C++ library for linear algebra. Journal of Open Source Software, 1(2): 26:1–26:2, June 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00026.

Sanderson:2017:PGP

[SC17] Conrad Sanderson and Ryan

Curtin. gmm diag and gmm full:

C++ classes for multi-threaded Gaussian mixture models and expectation-maximisation. Journal of Open Source Software, 2(18):365:1–365:2, October 2017. CODEN ????

ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00365.

Scott:2017:PSI

[Sco17]  Camille Scott.  shmlast:

An improved implementation of conditional reciprocal best hits with LAST and Python. Journal of Open Source Software, 2(9):142:1–142:4, January 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00142.

Sousa:2018:PBR

[SD18] David N. Sousa and João R. Daniel. bor: An R package for transforming behavioral observation records into data matrices. Journal of Open Source Software, 3(28):909:1–909:4, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00909.

Smith:2018:POP

[SG18] Daniel G. A. Smith and John-

                                        nie Gray.  opt einsum —

a Python package for optimizing contraction order for einsum-like expressions. Journal of Open Source Software, 3(26):753:1–753:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00753.

Suess:2017:PMM

[SH17] Daniel Suess and Milan Holzäpfel.

mpnum: A matrix product representation library for Python. Journal of Open Source Software, 2(20):465:1–465:2, December 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00465.

Sparks:2017:PGG

[SHN17] Adam H. Sparks, Tomislav Hengl, and Andrew Nelson. GSODR: Global summary daily weather data in R. Journal of Open Source Software, 2(10):177:1–177:2, February 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00177.

Siu:2017:PRP

[Siu17] Chapman Siu. Ramble: Parser combinator for R. Journal of Open Source Software, 2 (11):209:1, March 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00209.

Schep:2017:PII

[SK17] Alicia N. Schep and Sarah K.

Kummerfeld. iheatmapr: Interactive complex heatmaps in R. Journal of Open Source Software, 2(16):359:1, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00359.

Stegmaier:2017:PGR

[SKW17] Philip Stegmaier, Alexander Kel, and Edgar Wingender. geneXplainR: An R interface for the geneXplain platform. Journal of Open Source Software, 2(18):412:1–412:3, October 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00412.

Sherman:2018:PBA

[SM18] Marcus D. Sherman and Ryan E. Mills. BAMnostic: an OS-agnostic toolkit for genomic sequence analysis. Journal of Open Source Software, 3(28):826:1–826:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00826.

Smith:2017:PUP

[Smi17] Andrew P. Smith. UKCensusAPI: python and R interfaces to the nomisweb UK census data API. Journal of Open Source Software, 2(19):408:1, November 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00408.

Smith:2018:PHC

[Smi18] Andrew P. Smith. humanleague: a C++ microsynthesis package with R and Python interfaces. Journal of Open Source Software, 3(25):629:1, May 2018. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00629.

Sheffield:2018:PSR

[SNR18] Nathan Sheffield, V. P. Nagraj, and Vince Reuter. simpleCache: R caching for reproducible, distributed, largescale projects. Journal of Open Source Software, 3(21):463:1, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00463.

Sochat:2017:SRO

[Soc17] Vanessa Sochat. Singularity registry: Open source registry for singularity images. Journal of Open Source Software, 2(18):426:1–426:3, October 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00426.

Sochat:2018:EFR

[Soc18a] Vanessa Sochat. The experiment factory: Reproducible experiment containers. Journal of Open Source Software, 3(22):521:1–521:4, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00521.

Sochat:2018:PCO

[Soc18b] Vanessa Sochat. Containershare:

Open Source Registry to build, test, deploy with CircleCI.

Journal of Open Source Software, 3(28):878:1–878:3, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00878.

Sochat:2018:PHC

[Soc18c]  Vanessa Sochat.  HelpMe

command line helper utility. Journal of Open Source Software, 3(26):775:1–775:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00775.

Sommer:2017:PPI

[Som17]  Philipp S. Sommer.  The

psyplot interactive visualization framework. Journal of Open Source Software, 2(16): 363:1, August 2017. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00363.

Seshadri:2017:PEQ

[SP17] Pranay Seshadri and Geoffrey

Parks. Effective-Quadratures

(EQ): Polynomials for computational engineering studies. Journal of Open Source Software, 2(11):166:1–166:2, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00166.

Sparks:2017:PGU

[Spa17] Adam H. Sparks. getCRUCLdata: Use and explore CRU CL v.

2.0 climatology elements in R. Journal of Open Source Software, 2(12):230:1, April 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00230.

Spielman:2018:PPP

[Spi18] Stephanie J. Spielman. phyphy:

Python package for facilitating the execution and parsing of HyPhy standard analyses. Journal of Open Source Software, 3(21):514:1, January 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00514.

Sparks:2017:PBF

[SPPP17] Adam H. Sparks, Mark Padgham, Hugh Parsonage, and Keith Pembleton. bomrang: Fetch Australian Government Bureau of Meteorology data in R. Journal of Open Source Software, 2(17):411:1–411:2, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00411.

Shi:2017:CCD

[SPSH^{+}17] Sinan Shi, David Pérez-Suárez, Steve Harris, Niall MacCallum, David Brealey, Mervyn Singer, and James Hetherington. Critical care data processing tools. Journal of Open Source Software, 2(20):513:1–513:4, December 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00513.

Silge:2016:PTT

[SR16] Julia Silge and David Robinson. tidytext: Text mining and analysis using tidy data principles in R. Journal of Open Source Software, 1 (3):37:1–37:3, July 2016. CO-DEN ???? ISSN 2475-9066.

                                        URL  http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00037.

Smith:2018:PUU

[SR18] Andrew P. Smith and Tom Russell. ukpopulation: unified national and subnational population estimates and projections, including variants. Journal of Open Source Software, 3(28):803:1–803:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00803.

Schlegel:2018:PHC

[SS18] Robert W. Schlegel and Albertus J. Smit. heatwaveR: A central algorithm for the detection of heatwaves and cold-spells. Journal of Open Source Software, 3(27):821:1– 821:2, July 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00821.

Stadler:2017:CCP

[Sta17]  Konstantin Stadler.  The

country converter coco — a Python package for converting country names between different classification schemes. Journal of Open Source Software, 2(16):332:1–332:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00332.

Salmon:2017:PRR

[SVM^{+}17] Maëlle Salmon, Sreekanth Vakacherla, Carles Milà, Julian D. Marshall, and Cathryn

Tonne. rtimicropem: an R

package supporting the analysis of RTI MicroPEM output files. Journal of Open Source Software, 2(16):333:1, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00333.

Tang:2018:PAR

[Tan18] Yuan Tang. autoplotly: An R package for automatic generation of interactive visualizations for statistical results. Journal of Open Source Software, 3(24):657:1–657:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00657.

Tauber:2016:PPP

[Tau16] J. K. Tauber. pyuca: a Python implementation of the Unicode Collation Algorithm. Journal of Open Source Software, 1(1): 21:1, May 2016. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00021.

Taylor:2018:PPP

[Tay18] Shawn David Taylor. pyPhenology: A python framework for plant phenology modelling. Journal of Open Source Software, 3(28):827:1–827:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00827.

Topping:2018:PPA

[TCR18] David Topping, Paul Connolly, and Jonathan Reid. PyBox: An automated box-model generator for atmospheric chemistry and aerosol simulations. Journal of Open Source Software, 3(28):755:1–755:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00755.

Thorley:2018:PCR

[Tho18] Joe Thorley. checkr: An R package for assertive programming. Journal of Open Source Software, 3(23):624:1, March 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00624.

Tierney:2017:PVV

[Tie17] Nicholas Tierney. visdat: Visualising whole data frames. Journal of Open Source Software, 2(16):355:1–355:2, August 2017. CODEN ????

ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00355.

Tocknell:2018:PHT

[Toc18] James Tocknell. h5preserve:

Thin wrapper around h5py, inspired by camel. Journal of Open Source Software, 3(22):581:1–581:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00581.

Torfi:2018:PSL

[Tor18] Amirsina Torfi. SpeechPy — a library for speech processing and recognition. Journal of Open Source Software, 3(27):749:1–749:5, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00749.

Tsyganov:2018:PRP

[TPAP18] Kirill Tsyganov, Andrew James Perry, Stuart Kenneth Archer, and David Powell. RNAsik: A pipeline for complete and reproducible RNA-seq analysis that runs anywhere with speed and ease. Journal of Open Source Software, 3(28): 583:1–583:3, August 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00583.

Tushar:2017:PFI

[TR17] Abhinav Tushar and Nicholas G.

Reich. flusight: interactive visualizations for infectious disease forecasts. Journal of Open Source Software, 2(13):231:1– 231:2, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00231.

Terhoeven:2018:PRG

[TSH18] Niklas Terhoeven, Jörg Schultz, and Thomas Hackl. reper: Genome-wide identification, classification and quantification of repetitive elements without an assembled genome. Journal of Open Source Software, 3(22):527:1–527:4, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00527.

Theodoropoulos:2016:PHT

[TSS16] Christos Theodoropoulos, Nikolaos Skoulikidis, and Anastasios Stamou. Habfuzz: A tool to calculate the instream hydraulic habitat suitability using fuzzy logic and fuzzy Bayesian inference. Journal of Open Source Software, 1(6):82:1–82:2, October 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00082.

Turner:2018:PQR

[Tur18] Stephen D. Turner. qqman: an R package for visualizing GWAS results using Q-Q and Manhattan plots. Journal of Open Source Software, 3(25):731:1–731:2, May 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00731.

VanderPlas:2016:PMC

[Van16] Jake VanderPlas. mst clustering:

Clustering via Euclidean minimum spanning trees. Journal of Open Source Software, 1(1): 12:1, May 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00012.

Vantas:2018:PHR

[Van18] Konstantinos Vantas. hydroscoper:

R interface to the Greek National Data Bank for Hydrological and Meteorological Information. Journal of Open Source Software, 3(23):625:1– 625:2, March 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00625.

VanHorn:2016:PMY

[VB16] Nicholas M. Van Horn and Aaron Beveridge. MassMine: Your access to data. Journal of Open Source Software, 1(8):50:1, December 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00050.

Vrbancic:2018:PNP

[VBM^{+}18] Grega Vrbančič, Lucija Brezočnik, Uroš Mlakar, Dušan Fister, and Iztok Fister, Jr. NiaPy: Python microframework for building nature-inspired algorithms. Journal of Open Source Software, 3(23):613:1–613:3, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00613.

Versypt:2018:PBO

[VCD18] Ashlee N. Ford Versypt, James D. Crall, and Biswadip Dey. BeeNestABM: An opensource agent-based model of spatiotemporal distribution of bumblebees in nests. Journal of Open Source Software, 3(27):718:1–718:2, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00718.

vanderHam:2018:PSD

[vdH18] Ruud van der Ham. salabim: discrete event simulation and animation in Python. Journal of Open Source Software, 3(27):767:1–767:2, July 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00767.

vanHeeringen:2017:PGD

[vH17a] Simon J. van Heeringen. genomepy: download genomes the easy way. Journal of Open Source Software, 2(16):320:1, August 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00320.

vanHulten:2017:PCC

[vH17b] Marco van Hulten. ComPlot:

Comparison plotter to visually evaluate ocean model simulations. Journal of Open Source Software, 2(17):368:1–368:4, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00368.

Versypt:2017:PAO

[VHM17] Ashlee N. Ford Versypt, Grace K. Harrell, and Alexandra N. McPeak. ACEInhibPKPD: An open-source MATLAB app for a pharmacokinetic/ pharmacodynamic model of ACE inhibition. Journal of Open Source Software, 2 (17):340:1–340:2, September 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00340.

Vitolo:2017:PHH

[Vit17] Claudia Vitolo. hddtools: Hydrological data discovery tools. Journal of Open Source Software, 2(9):56:1–56:2, January 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00056.

Vitolo:2016:PRI

[VRT16] Claudia Vitolo, Andrew Russell, and Allan Tucker. rdefra:

Interact with the UK AIR Pollution Database from DEFRA. Journal of Open Source Software, 1(4):51:1–51:2, August 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00051.

Vitolo:2016:PFR

[VWDB16] Claudia Vitolo, Peter Wells, Martin Dobias, and Wouter Buytaert. fuse: An R package for ensemble hydrological modelling. Journal of Open Source Software, 1(8):52:1–52:2, December 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00052.

Waggoner:2018:PHP

[Wag18] Philip D. Waggoner. The hhi package: Streamlined calculation and visualization of Herfindahl–Hirschman index scores. Journal of Open Source Software, 3(28):828:1–828:5, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00828.

Wakefield:2016:PSI

[Wak16a] Matthew J. Wakefield. SurvivalVolume: interactive volume threshold survival graphs. Journal of Open Source Software, 1(8): 111:1–111:2, December 2016. CODEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00111.

Wakefield:2016:PXM

[Wak16b] Matthew J. Wakefield. Xenomapper:

Mapping reads in a mixed species context. Journal of Open Source Software, 1(1): 18:1, May 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00018.

Warren:2016:PRP

[War16]  Rene L. Warren.  RAILS

and Cobbler: Scaffolding and automated finishing of draft genomes using long DNA sequences. Journal of Open Source Software, 1(7):116:1– 116:2, November 2016. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00116.

Warren:2018:VGS

[War18]  René L. Warren.  Visual-

izing genome synteny with xmatchview. Journal of Open Source Software, 3(21):497:1– 497:4, January 2018. CO-DEN ???? ISSN 2475-9066.

URL  http://joss.theoj.
org/papers/10.21105/joss.
00497.

Wahls:2018:PFS

[WCP18] Sander Wahls, Shrinivas Chimmalgi, and Peter J. Prins. FNFT: A software library for computing nonlinear Fourier transforms. Journal of Open Source Software, 3(23):597:1–597:2, March 2018. CODEN ????

ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00597.

Wilkinson:2018:PPR

[WD18] Shaun P. Wilkinson and Simon K. Davy. phylogram: an R package for phylogenetic analysis with nested lists. Journal of Open Source Software, 3(26):790:1–790:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00790.

Wilson:2017:PLS

[WFA^{+}17] Lucas A. Wilson, John M. Fonner, Jason Allison, Oscar Esteban, Harry Kenya, and Marshall Lerner. Launcher: A simple tool for executing high throughput computing workloads. Journal of Open Source Software, 2(16):289:1–289:2, August 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00289.

Whalley:2018:PEE

[Wha18] Lucy D. Whalley. effmass: An effective mass package. Journal of Open Source Software, 3(28):797:1–797:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00797.

Willner:2017:PPP

[WHG17] Sven N. Willner, Corinne Hartin, and Robert Gieseke. pyhector: A Python interface for the simple climate model Hector. Journal of Open Source Software, 2(12):248:1–248:2, April 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00248.

Weisenburger:2017:PAA

[WHLM17] Kolby L. Weisenburger, Joseph Huehnerhoff, Emily M. Levesque, and Philip Massey. acronym: An automatic reduction pipeline for astronomical images. Journal of Open Source Software, 2(13):102:1–102:2, May 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00102.

Wiley:2018:PRR

[Wil18a]  David Wiley.  RISE: An

R package for RISE analysis. Journal of Open Source Software, 3(28):846:1, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00846.

Williams:2018:CPN

[Wil18b] Benjamin Williams. Combining a probability and a non-probability sample in a capture–recapture setting. Journal of Open Source Software, 3(28):886:1–886:2, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00886.

Wojdyr:2017:PUW

[Woj17] Marcin Wojdyr. UglyMol: a WebGL macromolecular viewer focused on the electron density. Journal of Open Source Software, 2(18):350:1–350:2, October 2017. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00350.

Wood:2018:PQQ

[Woo18]  Tobias C. Wood.  QUIT:

QUantitative imaging tools. Journal of Open Source Software, 3(26):656:1–656:2, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00656.

Wette:2018:POL

[WPK^{+}18] Karl Wette, Reinhard Prix, David Keitel, Matthew Pitkin, Christoph Dreissigacker, John T. Whelan, and Paola Leaci. OctApps: a library of Octave functions for continuous gravitational-wave data analysis. Journal of Open Source Software, 3(26):707:1–707:3, June 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00707.

Wrobel:2018:PRR

[Wro18] Julia Wrobel. register: Registration for exponential family functional data. Journal of Open Source Software, 3(22):557:1–557:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00557.

Wadhwa:2018:PTR

[WWDS18] Raoul R. Wadhwa, Drew F. K.

Williamson, Andrew Dhawan, and Jacob G. Scott. TDAstats: R pipeline for computing persistent homology in topological data analysis. Journal of Open Source Software, 3(28):860:1–860:3, August 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00860.

Willcox:2018:PPI

[WZ18] Donald E. Willcox and Michael Zingale. pynucastro: an interface to nuclear reaction rates and code generator for reaction network equations. Journal of Open Source Software, 3(23):588:1–588:3, March 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00588.

Yao:2017:PWM

[YK17] Andy Yu Zhu Yao and David Kane. walkr: MCMC sampling from non-negative convex polytopes. Journal of Open Source Software, 2(11):61:1, March 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00061.

Youngflesh:2018:PMT

[You18] Casey Youngflesh. MCMCvis: Tools to visualize, manipulate, and summarize MCMC output. Journal of Open Source Software, 3(24):640:1–640:3, April 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00640.

Yu:2018:PEU

[YS18] Chi-Lin Yu and Ching-Fan Sheu. EFAshiny: An userfriendly shiny application for exploratory factor analysis. Journal of Open Source Software, 3(22):567:1–567:2, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00567.

Zaghdoudi:2018:PIR

[Zag18] Taha Zaghdoudi. ivporbit: An R package to estimate the instrumental variables probit model. Journal of Open Source Software, 3(22):523:1, February 2018. CODEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/10.
                                        21105/joss.00523.

Zhang:2017:PIT

[ZB17] Yuxuan Zhang and Jean Bil-

                                        heux. ImagingReso: A tool

for neutron resonance imaging. Journal of Open Source Software, 2(19):407:1–407:3, November 2017. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00407.

Zeki:2018:PEE

[Zek18] Sebastian S. Zeki. EndoMineR for the extraction of endoscopic and associated pathology data from medical reports. Journal of Open Source Software, 3(24):701:1–701:2, April 2018. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00701.

Zhang:2016:ASG

[ZKM^{+}16] Zhao Zhang, Daniel S. Katz, Andre Merzky, Matteo Turilli, Shantenu Jha, and Yadu Nand. Application skeleton: Generating synthetic applications for infrastructure research. Journal of Open Source Software, 1(1):17:1–17:2, May 2016. CODEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/10.
21105/joss.00017.